### v2

# Extracting the expression for the field equations of a diffeomorphism invariant theory of gravity from surface term

Abstract
As a contribution towards the understanding for the field equations of diffeomorphism invariant theories of pure gravity, we demonstrate in great detail that the expression for the field equations of such theories can be derived within the perspective of the surface term coming from the variation of the Lagrangian. Specifically, starting with the surface term, we extract a symmetric rank-two tensor together with an anti-symmetric one out of this term with the variation operator replaced with the Lie derivative along an arbitrary vector field. By utilizing an equality stemming from the Lie derivative of the Lagrangian density along an arbitrary vector field, it is proved that the resulting symmetric rank-two tensor is identified with the functional derivative of the Lagrangian density with respect to the metric. Such a result further brings forth the expression for the field equations constructed from the symmetric rank-two tensor, which naturally rules out the derivative of the Lagrangian density with respect to the metric and coincides with the one for the Euler-Lagrange equations of motion. Furthermore, it is illustrated that the construction of the expression for the field equations from the symmetric rank-two tensor must be feasible as long as the variation operator in the variation equation of the Lagrangian is allowed to be substituted by the Lie derivative along an arbitrary vector field. On the other hand, as a byproduct, the anti-symmetric rank-two tensor turns out to be the Noether charge two-form. Subsequently, it is illustrated in detail that there exist two equivalent ways to identify the field equations arising from the surface term with the Euler-Lagrange equations. In addition, utilizing the surface term, together with two equalities derived from the variation equations of the Lagrangian with the variation operator substituted by the Lie derivative, we obtain some equivalence relationships associated to the conserved current and the expression for the field equations. Our results offer a straightforward support on the proposal in our previous work that the surface term gives a unified description for field equations and Noether charges in the context of theories of gravity admitting diffeomorphism invariance symmetry. What is more, we observe that the expression for the field equations can appear as a consequence for the vanishing of its divergence or the divergence for the off-shell Noether current.
1 Introduction
Given a Lagrangian based theory of gravity, the conventional approach to derive the gravitational field equations associated to this theory is to vary the Lagrangian with respect to the metric under the guidance of the variational principle, which is a powerful tool in field theories [1, 2, 3, 4, 5]. From a general perspective, the resulting equations of motion, referred to as the Euler-Lagrange equations in the literature as well, are expressed through the functional derivative of the Lagrangian with respect to the metric. This expression ordinarily gets involved in the derivative of the Lagrangian density with respect to the metric. Thus, the elimination of such a term renders it of possibility to simplify the equations of motion.
Within the paper [6], by utilizing an identity establishing the relation between the derivatives of the Lagrangian density with respect to the metric and the Riemann curvature tensor respectively, which appears as a consequence of the comparison between the results for the Lie derivative of the Lagrangian density evaluated in two different ways, Padmanabhan presented a compact form for the field equations of gravity without the derivative of the Lagrangian density with respect to the metric in the context of generally covariant metric theories. However, since the Lagrangians under consideration merely depend upon the metric and the Riemann curvature tensor, the application range of the resulting field equations is restricted to gravity theories equipped with the Lagrangian that is free of the covariant derivative(s) of the Riemann tensor. Subsequently, as a generalization to the higher derivative theories of gravity constructed out of the first-order covariant derivative of the Riemann tensor in addition to the metric and the Riemann tensor, Padmanabhan’s method was extended to derive the equations of motion for such theories in the paper [7]. Another generalization to the gravity theories in the presence of a vector field together with its first-order covariant derivative was given by the paper [8] quite recently. Till now, owing to the conciseness of the field equations presented by [6], they have gained a broad range of applications. Some of them can be found in the works [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
What is more, for a generally covariant Lagrangian-based metric theory of gravity with or without the covariant derivatives of the Riemann tensor, the variation of the Lagrangian with regard to the metric brings forth a surface term that naturally eliminates the derivative of the Lagrangian density with respect to the metric. As is what has been illustrated in our earlier work [20], such a surface term, which is also referred to as the symplectic potential current density [21], can be employed to generate the expression for the field equations of the gravitational field. Not unnaturally, throughout the whole process for doing so, one is able to keep away from any calculation involving the derivative of the Lagrangian density with respect to the metric. By contrast, in the framework of Padmanabhan’s method, it is unavoidable to deal with the identity associated to this quantity as it plays a prominent role in simplifying the equations of motion fully arising from the variation of the Lagrangian. Apparently, the field equations from the perspective of the surface term naturally inherit the feature for the surface term, that is, the derivative of the Lagrangian density with respect to the metric is completely excluded from them.
According to the approach put forward in Ref. [20], the surface term stemming from the variation of the Lagrangian plays a central role. Under the substitution for the variation operator by the Lie derivative along an arbitrary vector field, the resulting surface term can be decomposed into the linear combination of two components. One of them is the contraction between the vector field and a symmetric rank-two tensor, while the other consists of the covariant divergence of another anti-symmetric rank-two tensor. It can be verified that the symmetric rank-two tensor corresponds to the expression for the field equations by making use of an identity linking its divergence with that of an off-shell Noether current. Actually, the approach to produce the field equations based on the surface term provides a natural framework for unifying both the field equations and the conserved current. Apart from the extraction of the field equations, the Noether charge two-form, whose Hodge dual is referred to as Noether charge (D2)(D-2)-form in $D$-dimensional spacetimes within the papers [1, 22], is able to be produced simultaneously. In practice, it is just the anti-symmetric rank-two tensor encompassed by the surface term. As an application, this surface term based approach has been extended to the higher derivative theories of gravity endowed with Lagrangians built from the metric and the Riemann tensor, together with $i$th ($i=1,2,\cdot\cdot\cdot$) powers of the Beltrami-d’Alembertian operator acting on the latter [23]. In addition, utilizing the field equations presented by [20], the authors investigated the stellar axisymmetry theorem in the context of a broad class of metric theories of gravity admitting diffeomorphism invariance within the paper [24].
The present paper is an extension and development to our earlier work [20]. It mainly focuses on some fundamental problems on how to extract the expression for the field equations out of the surface term arising from the variation of the Lagrangian, such as whether the surface term must accommodate the rank-two symmetric tensor that is able to participate in the construction of the field equations, or what constitute the foundation for the derivation of the field equations from the surface term, or what characterize the expression for the equations of motion extracted out of the surface term. Regarding these issues, this paper encompasses four main objectives. First, building upon the work [20], we attempt to seek another route to confirming that the expression of the field equations for gravity theories admitting diffeomorphism invariance symmetry can be extracted out of the surface term. On the basis of the two routes, our intention is to establish the sufficient and necessary conditions for that the surface term is able to produce such an expression. Thus it will be verified that deriving field equations from the surface term perspective indeed establishes on a solid foundation. Due to this, we expect that our approach will enhance the understanding of the implications for utilizing the symmetries of the Lagrangian to yield the equations of motion. Second, due to that the surface term is able to provide a unified description towards the conserved Noether current and the expression for the field equations, we anticipate giving some insights into the relationships between both of them. Furthermore, inspired with the results, we plan to verify that the expression for the equations of motion can stem from the disappearance of its covariant divergence. Third, in terms of the surface term perspective towards the derivation for the expression of the field equations, we observe that three equations (see Eqs. (25), (39) and (52) below for them), which result from performing the replacement of the variation operator with the Lie derivative towards the surface term and the variation equations of the Lagrangian, occupy a decisive position in producing the desired expression for the equations of motion. Thus, we plan to reveal the deep connections among them. Fourth, within the context of the diffeomorphism invariant theories of gravity comprising arbitrary order covariant derivatives of the Riemann tensor, under the guidance of the expression for the field equations within the surface term perspective, we attempt to fill the gap in the derivation for the field equations of these theories by strictly following Padmanabhan’s method. At the same time, we will carry out a detailed comparison between the two approaches in order to offer an improved understanding of Padmanabhan’s method. In addition, we intend to obtain the explicit Noether charge two-forms for these theories.
The present paper is structured as follows. Within Section 2, the equations of motion and the surface term for diffeomorphism invariant theories of pure gravity involving the covariant derivatives of the Riemann tensor are figured out via the variation of the Lagrangians for these theories. Within Section 3, under the replacement for the variation operator by the Lie derivative along an arbitrary smooth vector field, the surface term is put into the form consisting of a symmetric rank-two tensor and an anti-symmetric one. By using an equality associated to the partial derivative of the Lagrangian density with respect to the metric, which is derived out of the Lie derivative of the Lagrangian density, then it is demonstrated that the symmetric rank-two tensor gives rise to the field equations and the anti-symmetric rank-two tensor represents the Noether charge two-form corresponding to the Noether current. Within Section 4, we make a detailed comparison between two different ways for identifying the expression for the Euler-Lagrange equations of motion with the one arising from the surface term. By the way, we obtain some equivalence relations relative to the conserved current and the expression for the field equations. Within Section 5, we reveal several relationships among three significant equations arising from the replacement of the variation operator with the Lie derivative along an arbitrary vector field in the surface term and the variation equations of the Lagrangian. Along the way, we explore the sufficient and necessary condition for the derivation of the field equations from the surface term. Within Section 6, through straightforward calculations, it is proved that the expression for the field equations originating from the surface term is divergence-free. Within Section 7, in terms of all the aforementioned analysis, the procedure for the derivation of the field equations from the surface term perspective is put forward. According to the procedure, we derive the explicit expressions for the surface term and the field equations associated to the generic theories of gravity endowed with the Lagrangian relying on the metric, the Riemann tensor and the covariant derivative operator. This section encompasses the main idea of the present work. The last section contains our conclusions and discussions.
2 Euler-Lagrange equations and surface terms resulting from the variation of the Lagrangian
In this section, we are going to derive field equations and surface terms by means of varying the Lagrangian with regard to the metric tensor.
As a beginning, we take into account a generic Lagrangian in arbitrary dimensions that is invariant under a change of coordinates on the spacetime manifold. Specifically, this diffeomorphism invariant Lagrangian is supposed to be built out of the metric together with $i$th-order $(i=0,1,\cdot\cdot\cdot,m)$ covariant derivatives of the Riemann curvature tensor. It has the form [1]
$\mathcal{L}=\sqrt{-g}L\left(g^{\mu\nu},R_{\mu\nu\rho\sigma},\nabla_{\lambda_{1}}R_{\mu\nu\rho\sigma},\cdot\cdot\cdot,\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{m}}R_{\mu\nu\rho\sigma}\right)\,.$ (1)
Computing straightforwardly the variation of the Lagrangian density $L$ with respect to all of its constituent variables, that is, the inverse metric $g^{\mu\nu}$ and the covariant derivatives of the Riemann curvature tensor $\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}$s, we obtain the result
$\delta{L}=\frac{\partial{L}}{\partial{g}^{\mu\nu}}\delta{g}^{\mu\nu}+\sum^{m}_{i=0}Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}\,$, (2)
where the rank-$(i+4)$ tensor $Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$, standing for the derivative of the Lagrangian density with respect to the variable $\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}$, is presented by
$Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}=Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\gamma\lambda}\Delta^{\mu\nu\rho\sigma}_{\alpha\beta\gamma\lambda}=\frac{\partial{L}}{\partial\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}}\,.$ (3)
The $\Delta^{\mu\nu\rho\sigma}_{\alpha\beta\gamma\lambda}$ tensor in Eq. (3), employed to characterize the same index symmetries as those for the Riemann tensor, is defined in terms of the Kronecker delta through
[…]
Within Eqs. (2) and (3), it is to be understood that no covariant derivatives act on the Riemann tensor when $i=0$ and $Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$ with $i=0$ represents $Z_{(0)}^{\mu\nu\rho\sigma}=\frac{\partial{L}}{\partial{R}_{\mu\nu\rho\sigma}}$. In terms of Eq. (2), the variation of the Lagrangian (1) is written down as
$\delta\mathcal{L}=\sqrt{-g}\left(\delta{L}-\frac{1}{2}{L}{g}_{\mu\nu}\delta{g}^{\mu\nu}\right)\,.$ (5)
For convenience, throughout this work, we introduce a third-rank tensor ${U}^{\mu\alpha\beta}_{(i)}$ defined in terms of $Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$s through
${U}^{\mu\alpha\beta}_{(i)}$ $=$ $4\sum^{i}_{k=1}(-1)^{k}\Big(\nabla_{\lambda_{k-1}}\cdot\cdot\cdot\nabla_{\lambda_{1}}Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{k-1}\alpha\lambda_{k+1}\cdot\cdot\cdot\lambda_{i}\beta\gamma\rho\sigma}\Big)\nabla_{\lambda_{k+1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R^{\mu}_{~\gamma\rho\sigma}$ (6)
$+\sum^{i-1}_{k=1}\sum^{i}_{\ell=k+1}(-1)^{k}\Big(\nabla_{\lambda_{k-1}}\cdot\cdot\cdot\nabla_{\lambda_{1}}Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{k-1}\alpha\lambda_{k+1}\cdot\cdot\cdot\lambda_{\ell-1}\beta\lambda_{\ell+1}\cdot\cdot\cdot\lambda_{i}\gamma\lambda\rho\sigma}\Big)\times$
$\times\nabla_{\lambda_{k+1}}\cdot\cdot\cdot\nabla_{\lambda_{\ell-1}}\nabla^{\mu}\nabla_{\lambda_{\ell+1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\gamma\lambda\rho\sigma}\,.$
For instance, in the above equation, ${U}^{\mu\alpha\beta}_{(0)}=0$ and ${U}^{\mu\alpha\beta}_{(1)}=-4Z_{(1)}^{\alpha\beta\gamma\rho\sigma}R^{\mu}_{~\gamma\rho\sigma}$. By the aid of the third-rank tensor ${U}^{\mu\alpha\beta}_{(i)}$, the scalar $Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}$ can be put into the form
$Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ $=$ $(-1)^{i}\big(\nabla_{\lambda_{i}}\cdot\cdot\cdot\nabla_{\lambda_{1}}Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\big)\delta{R}_{\mu\nu\rho\sigma}$ (7)
$-{B}^{\mu\nu}_{(i)}\delta{g}_{\mu\nu}+\nabla_{\mu}\bar{\Theta}^{\mu}_{(i)}\,.$
Within Eq. (7), the second-rank symmetric tensor ${B}^{\mu\nu}_{(i)}$ is defined in terms of the divergences of the tensor ${U}^{\mu\alpha\beta}_{(i)}$ as
${B}^{\mu\nu}_{(i)}=\frac{1}{2}\nabla_{\lambda}\left({U}^{(\mu\nu)\lambda}_{(i)}-{U}^{\lambda(\mu\nu)}_{(i)}+{U}^{(\mu|\lambda|\nu)}_{(i)}\right)\,$, (8)
while the vector $\bar{\Theta}^{\mu}_{(i)}$, appearing as a result of peeling off the covariant derivative operator in $\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$, has the form
$\bar{\Theta}^{\mu}_{(i)}$ $=$ $\sum^{i}_{k=1}(-1)^{k-1}\Big(\nabla_{\lambda_{k-1}}\cdot\cdot\cdot\nabla_{\lambda_{1}}{Z}^{\lambda_{1}\cdot\cdot\cdot\lambda_{k-1}\mu\lambda_{k+1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}_{(i)}\Big)\delta\nabla_{\lambda_{k+1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\alpha\beta\rho\sigma}$ (9)
$+\frac{1}{2}\left({U}^{(\alpha|\mu|\beta)}_{(i)}+{U}^{(\alpha\beta)\mu}_{(i)}-{U}^{\mu(\alpha\beta)}_{(i)}\right)\delta{g}_{\alpha\beta}\,.$
Particularly, $\bar{\Theta}^{\mu}_{(0)}=0$. Obviously, the dependence of $\bar{\Theta}^{\mu}_{(i)}$ on $\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ and $\delta{g}_{\alpha\beta}$ is linear. As a consequence of Eq. (7), equation (5) is recast into
$\delta\mathcal{L}$ $=$ $\sqrt{-g}\left[\left(\frac{\partial{L}}{\partial{g}^{\mu\nu}}-\frac{1}{2}{L}{g}_{\mu\nu}+{B}_{\mu\nu}\right)\delta g^{\mu\nu}+{P}^{\mu\nu\rho\sigma}\delta{R}_{\mu\nu\rho\sigma}\right]$ (10)
[…]
In Eq. (10), the second-rank symmetric tensor ${B}^{\mu\nu}$ is defined as
${B}^{\mu\nu}=\sum^{m}_{i=0}{B}^{\mu\nu}_{(i)}\,$, (11)
and the rank-four tensor ${P}^{\mu\nu\rho\sigma}$ is in the form
${P}^{\mu\nu\rho\sigma}=\sum^{m}_{i=0}(-1)^{i}\nabla_{\lambda_{i}}\cdot\cdot\cdot\nabla_{\lambda_{1}}Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\,.$ (12)
Furthermore, after utilizing [20]
${P}^{\mu\nu\rho\sigma}\delta{R}_{\mu\nu\rho\sigma}=-\big({P}_{\mu}^{~\lambda\rho\sigma}R_{\nu\lambda\rho\sigma}+2\nabla^{\rho}\nabla^{\sigma}{P}_{\rho\mu\nu\sigma}\big)\delta{g}^{\mu\nu}+\nabla_{\mu}\Theta^{\mu}_{P}\,$, (13)
with the vector $\Theta^{\mu}_{P}$ given by
$\Theta^{\mu}_{P}=2{P}^{\mu\nu\rho\sigma}\nabla_{\sigma}\delta{g}_{\rho\nu}-2(\delta{g}_{\nu\rho})\nabla_{\sigma}{P}^{\mu\nu\rho\sigma}\,$, (14)
we find that equation (10) is ultimately written as the conventional form for the variation of the Lagrangian [1, 21, 22], namely,
$\delta\mathcal{L}=\sqrt{-g}\left({E}_{\mu\nu}\delta{g}^{\mu\nu}+\nabla_{\mu}\Theta^{\mu}\right)\,.$ (15)
Within the above equation, the expression ${E}_{\mu\nu}$ for equations of motion is read off as
${E}_{\mu\nu}$ $=$ $\frac{1}{\sqrt{-g}}\frac{\delta\mathcal{L}}{\delta{g}^{\mu\nu}}=\frac{\delta{L}}{\delta{g}^{\mu\nu}}-\frac{1}{2}{L}{g}_{\mu\nu}$ (16)
$=$ $\frac{\partial{L}}{\partial{g}^{\mu\nu}}-\frac{1}{2}{L}{g}_{\mu\nu}+{B}_{\mu\nu}-{P}_{\mu}^{~\lambda\rho\sigma}R_{\nu\lambda\rho\sigma}-2\nabla^{\rho}\nabla^{\sigma}{P}_{\rho\mu\nu\sigma}\,.$
In terms of Eq. (16), the field equations, that is, the Euler-Lagrange equations of motion, are written down as
${E}_{\mu\nu}=0\,.$ (17)
The surface term $\Theta^{\mu}$ in Eq. (15) incorporates $\Theta^{\mu}_{P}$ and $\bar{\Theta}^{\mu}_{(i)}$s in the way
$\Theta^{\mu}=\Theta^{\mu}_{P}+\sum^{m}_{i=0}\bar{\Theta}^{\mu}_{(i)}\,$, (18)
which fulfills Eq. (136). Apparently, here the surface term $\Theta^{\mu}$ merely arises from the $Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}$ terms within the variation equation (2) for the Lagrangian density and it naturally rules out the symmetric tensor $\frac{\partial{L}}{\partial{g}^{\mu\nu}}$ or $\frac{\partial{L}}{\partial{g}_{\mu\nu}}$. This accordingly causes the vanishing of such a tensor in the field equations resulting from the surface term as is what will be shown in the next section. Like usual, the expression (16) for the equations of motion is obtained by following the variational principle to directly vary the Lagrangian with respect to the metric. It adequately comes from the terms proportional to the variation of the metric and is irrelevant to the surface terms. This implies that the surface term can be ignored if the purpose is only to derive the Euler-Lagrange equations for gravitational field. However, it will be illustrated below that the surface term is of great importance throughout our analysis. It encodes all the information for the expressions of the field equations and the conserved Noether current. In addition, owing to the fact that both the second-rank tensors $\frac{\partial{L}}{\partial{g}^{\mu\nu}}$ and ${B}_{\mu\nu}$ are symmetric, together with the equality [6, 11]
${P}_{[\mu}^{~~\lambda\rho\sigma}R_{\nu]\lambda\rho\sigma}+2\nabla^{\rho}\nabla^{\sigma}{P}_{\rho[\mu\nu]\sigma}=0\,$, (19)
it is always guaranteed that ${E}_{\mu\nu}={E}_{\nu\mu}$.
3 Identifying the expression for the field equations arising from surface term with the Euler-Lagrange expression
In this section, we deal with the surface term under the transformation of the variation operator to the Lie derivative along an arbitrary smooth vector field. From the surface term via such a transformation, a rank-two symmetric tensor irrelevant to the vector is extracted. By means of establishing the relationship among the rank-two tensor $\frac{\partial{L}}{\partial{g}^{\mu\nu}}$ and the rank-$(i+4)$ ones $Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$s, then it is strictly proved that the resulting rank-two symmetric tensor is identified with the functional derivative of the Lagrangian density with respect to the metric. This equality further gives rise to that the field equations arising from the surface term are in agreement with the Euler-Lagrange equations. Apart from this, it is illustrated that the off-shell Noether current can be derived out of the surface term as well.
First, with the surface term (18) at hand, we concentrate on the behaviour of $\Theta^{\mu}$ under the transformation $\delta\rightarrow\mathcal{L}_{\zeta}$, where $\mathcal{L}_{\zeta}$ denotes the Lie derivative along an arbitrary smooth vector field $\zeta^{\mu}$. Our main goal is to put the surface term $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ into the form displayed by Eq. (101). The detailed procedure for doing so is presented in Appendix A. So we directly adopt the related results there. As a beginning, we perform the transformation $\delta\rightarrow\mathcal{L}_{\zeta}$ to the vector $\Theta^{\mu}_{P}$, which leads to [20]
$\Theta^{\mu}_{P}(\delta\rightarrow\mathcal{L}_{\zeta})=2\zeta_{\nu}\left({P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}-2\nabla_{\rho}\nabla_{\sigma}{P}^{\rho\mu\nu\sigma}\right)-\nabla_{\nu}{K}^{\mu\nu}_{P}\,$, (20)
in which the skew-symmetric tensor ${K}^{\mu\nu}_{P}$ is in the form
${K}^{\mu\nu}_{P}=2{P}^{\mu\nu\rho\sigma}\nabla_{\rho}\zeta_{\sigma}+4\zeta_{\rho}\nabla_{\sigma}{P}^{\mu\nu\rho\sigma}-6{P}^{\mu[\nu\rho\sigma]}\nabla_{\rho}\zeta_{\sigma}\,.$ (21)
Besides, as a straightforward application of Eq. (121), replacing the variation operator $\delta$ in $\bar{\Theta}^{\mu}_{(i)}$ with the Lie derivative $\mathcal{L}_{\zeta}$ gives rise to
$\bar{\Theta}^{\mu}_{(i)}(\delta\rightarrow\mathcal{L}_{\zeta})=2\zeta_{\nu}{W}^{\mu\nu}_{(i)}-\nabla_{\nu}\bar{K}^{\mu\nu}_{(i)}\,.$ (22)
[…]
$W^{\mu\nu}_{(i)}$ $=$ $\frac{1}{2}\sum^{i}_{k=1}(-1)^{k-1}\Big(\nabla_{\lambda_{k-1}}\cdot\cdot\cdot\nabla_{\lambda_{1}}Z^{\lambda_{1}\cdot\cdot\cdot\lambda_{k-1}\mu\lambda_{k+1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}_{(i)}\Big)\nabla^{\nu}\nabla_{\lambda_{k+1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\alpha\beta\rho\sigma}$ (23)
[…]
and the anti-symmetric rank-two tensor $\bar{K}^{\mu\nu}_{(i)}=-\zeta_{\lambda}\Psi^{\lambda\mu\nu}_{(i)}$ is presented by
[…]
Thus, by means of the combination of Eqs. (20) and (22), the surface term $\Theta^{\mu}$ under the transformation $\delta\rightarrow\mathcal{L}_{\zeta}$ turns into the form having the same structure as the one in Eq. (101), that is,
$\boxed{\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})=2\zeta_{\nu}{X}^{\mu\nu}-\nabla_{\nu}{Q}^{\mu\nu}}\,.$ (25)
In Eq. (25), the rank-two tensor ${X}^{\mu\nu}$ given by Eq. (126) is recast into
${X}^{\mu\nu}={P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}-2\nabla_{\rho}\nabla_{\sigma}{P}^{\rho\mu\nu\sigma}+{W}^{\mu\nu}\,$, (26)
with the ${W}^{\mu\nu}$ tensor given by
${W}^{\mu\nu}=\sum^{m}_{i=0}W^{\mu\nu}_{(i)}\,.$ (27)
In terms of the anti-symmetric tensors $\bar{K}^{\mu\nu}_{(i)}$s, the anti-symmetric tensor ${Q}^{\mu\nu}$ given by Eq. (127) can be rewritten as
${Q}^{\mu\nu}$ $=$ $2{P}^{\mu\nu\rho\sigma}\nabla_{\rho}\zeta_{\sigma}+4\zeta_{\rho}\nabla_{\sigma}{P}^{\mu\nu\rho\sigma}-6{P}^{\mu[\nu\rho\sigma]}\nabla_{\rho}\zeta_{\sigma}$ (28)
[…]
As what has been demonstrated in Appendix A, in order to guarantee that the counterpart of the surface term $\Theta^{\mu}$ under $\delta\rightarrow\mathcal{L}_{\zeta}$ takes on the structure displayed by Eq. (25), here we stress once again that the single requirement is that there exists an appropriate rank-two tensor ${X}^{\mu\nu}$ enabling the difference between $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ and $2\zeta_{\nu}{X}^{\mu\nu}$ to be divergence-free. In view of our straightforward calculations, the existence of ${X}^{\mu\nu}$ is always guaranteed within the framework of the diffeomorphism invariant gravity described by the Lagrangian (1). This is also supported by Eq. (64) or (141). What is more, within Appendix A, it is illustrated that the rank-two tensor contracted with the arbitrary vector $\zeta^{\mu}$ in the first term at the right hand side of Eq. (101) is unique under the condition that the surface term with the variation operator replaced by the Lie derivative exhibits the structure appearing in this equation. That is to say, the tensor ${X}^{\mu\nu}$ maintains unchanged accompanied by the alterations to the anti-symmetric tensor ${Q}^{\mu\nu}$, which is determined up to the divergence of an arbitrary three-form.
Next, we switch to focusing on the derivation of the field equations and conserved current out of the surface term. As what has been demonstrated in Ref. [20], the surface term (25) completely accounts for the equations of motion associated to the Lagrangian (1). In order to avoid confusion with the Euler-Lagrange expression ${E}^{\mu\nu}$ given by Eq. (16), throughout this paper, the second-rank tensor $\tilde{E}^{\mu\nu}$ is utilized to represent the expression for such equations of motion stemming from the surface term, and it is expressed in terms of the second-rank tensor ${X}^{\mu\nu}$ as
$\tilde{E}^{\mu\nu}={X}^{\mu\nu}-\frac{1}{2}{L}g^{\mu\nu}\,.$ (29)
According to the procedure put forward in Ref. [20], which can be also found in Appendix A, an off-shell Noether current involving both the expression for the field equations and the surface term with $\delta$ substituted by $\mathcal{L}_{\zeta}$ plays a fundamental role in producing the field equations under the condition that the surface term $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ possesses the form (25). The variation equation (15) of the Lagrangian under the transformation $\delta\rightarrow\mathcal{L}_{\zeta}$ generates an identity that connects the divergence of this current with that for the expression of the Euler-Lagrange equations (see Eq. (52) or (103) for it). In essence, this identity establishes a concrete connection between ${X}^{\mu\nu}$ and ${E}^{\mu\nu}$. Attributed to the fact that such an identity is valid for any vector field $\zeta^{\mu}$, there has to exist the one-to-one correspondence between the second-rank tensor ${X}^{\mu\nu}$ and the expression ${E}^{\mu\nu}$ for the Euler-Lagrange equations in such a way that ${E}^{\mu\nu}={X}^{\mu\nu}-\frac{1}{2}{L}g^{\mu\nu}$, as is shown by Eq. (107). By comparison with Eq. (29), then one immediately arrives at the conclusion $\tilde{E}^{\mu\nu}={E}^{\mu\nu}$. However, for the sake of offering a particularly straightforward proof rather than resorting to the identity related to the off-shell Noether current to verify that $\tilde{E}^{\mu\nu}$ indeed coincides with ${E}^{\mu\nu}$ like in Ref. [20], in what follows, we completely abandon the aforementioned procedure from the off-shell Noether current perspective. As an alternative route to doing so, we directly compare ${X}_{\mu\nu}$ with $\frac{\delta{L}}{\delta{g}^{\mu\nu}}$ to confirm that there exists no difference between both of them. This leads to $\tilde{E}^{\mu\nu}={E}^{\mu\nu}$ in a quite direct manner.
To achieve the goal, it is the time for us to turn to the difference $(\Delta{E})_{\mu\nu}$ between both the tensors $\frac{\delta{L}}{\delta{g}^{\mu\nu}}$ and ${X}_{\mu\nu}$, namely,
$(\Delta{E})_{\mu\nu}={X}_{\mu\nu}-\frac{\delta{L}}{\delta{g}^{\mu\nu}}\,.$ (30)
Alternatively, the rank-two tensor $(\Delta{E})_{\mu\nu}$ can be defined in terms of the difference between ${E}_{\mu\nu}$ and $\tilde{E}_{\mu\nu}$ as
$(\Delta{E})_{\mu\nu}$ $=$ ${E}_{\mu\nu}-\tilde{E}_{\mu\nu}$ (31)
$=$ $\frac{\partial{L}}{\partial{g}^{\mu\nu}}+{B}_{\mu\nu}-2{P}_{\mu}^{~\lambda\rho\sigma}{R}_{\nu\lambda\rho\sigma}-{W}_{\mu\nu}\,.$
The substitution of Eqs. (11) and (27) into Eq. (31) yields
[…]
$\times\nabla^{\nu}\nabla_{\lambda_{k+1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\alpha\beta\rho\sigma}$
$+{g}^{\mu\rho}{g}^{\nu\sigma}\frac{\partial{L}}{\partial{g}^{\rho\sigma}}-2{P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}+\frac{1}{2}\sum^{m}_{i=0}\nabla_{\lambda}{U}^{\nu\lambda\mu}_{(i)}\,$,
where the rank-two contravariant tensor $(\Delta{E})^{\mu\nu}={g}^{\mu\rho}{g}^{\nu\sigma}(\Delta{E})_{\rho\sigma}$. For the sake of simplifying Eq. (32), it is useful to introduce the following equality
[…]
$-Z^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\gamma\rho\sigma}_{(i)}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}^{\nu}_{~\gamma\rho\sigma}\,$,
where the third-rank tensor ${A}^{\mu}_{(i,n;\ell)}$ $(n=1,2,\cdot\cdot\cdot,i;\ell=0,1,\cdot\cdot\cdot,n-1)$ is defined as
${A}^{\nu\alpha\mu}_{(i,n;\ell)}=(-1)^{n-\ell}\Big(\nabla_{\lambda_{n-\ell-1}}\cdot\cdot\cdot\nabla_{\lambda_{1}}Z^{\lambda_{1}\cdot\cdot\cdot\lambda_{n-\ell-1}\alpha\lambda_{n-\ell+1}\cdot\cdot\cdot\lambda_{i}\mu\gamma\rho\sigma}_{(i)}\Big)\nabla_{\lambda_{n-\ell+1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}^{\nu}_{~\gamma\rho\sigma}\,$, (34)
with ${A}^{\nu\alpha\mu}_{(0,n;\ell)}\equiv 0$ and ${A}^{\nu\alpha\mu}_{(i,n;\ell)}={A}^{\nu\alpha\mu}_{(i,p;q)}$ under the condition $n-\ell=p-q$. Particularly, the contraction between the indices $\mu$ and $\nu$ in Eq. (33) produces the equality related to the vector ${A}^{\mu}_{(i,n;\ell)}$ in [25]. By the aid of equality (33), the rank-two tensor ${P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}$ can be written as
${P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}$ $=$ $\sum^{m}_{i=0}{Z}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\lambda\rho\sigma}_{(i)}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}^{\nu}_{~\lambda\rho\sigma}+\sum^{m}_{i=0}\sum^{i}_{k=1}\nabla_{\lambda}{A}^{\nu\lambda\mu}_{(i,i;i-k)}\,.$ (35)
What is more, it can be verified that the divergence for the third-rank tensor ${U}^{\nu\lambda\mu}_{(i)}$ is connected with the sum for the divergence of ${A}^{\nu\lambda\mu}_{(i,i;i-k)}$ in the following manner
$\nabla_{\lambda}{U}^{\nu\lambda\mu}_{(i)}$ $=$ $\sum^{i}_{k=1}(-1)^{k-1}\Big(\nabla_{\lambda_{k-1}}\cdot\cdot\cdot\nabla_{\lambda_{1}}Z^{\lambda_{1}\cdot\cdot\cdot\lambda_{k-1}\mu\lambda_{k+1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}_{(i)}\Big)\nabla^{\nu}\nabla_{\lambda_{k+1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\alpha\beta\rho\sigma}$ (36)
$-\sum^{i}_{k=1}{Z}^{\lambda_{1}\cdot\cdot\cdot\lambda_{k-1}\mu\lambda_{k+1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}_{(i)}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{k-1}}\nabla^{\nu}\nabla_{\lambda_{k+1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\alpha\beta\rho\sigma}$
$+4\sum^{i}_{k=1}\nabla_{\lambda}{A}^{\nu\lambda\mu}_{(i,i;i-k)}\,.$
Substituting Eqs. (35) and (36) into Eq. (32), one observes that the tensor $(\Delta{E})^{\mu\nu}$ is reformulated as the form in the absence of the arbitrary order covariant derivatives of the ${Z}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}_{(i)}$ tensor, that is,
$(\Delta{E})^{\mu\nu}$ $=$ $-\frac{1}{2}\sum^{m}_{i=0}\sum^{i}_{k=1}{Z}^{\lambda_{1}\cdot\cdot\cdot\lambda_{k-1}\mu\lambda_{k+1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}_{(i)}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{k-1}}\nabla^{\nu}\nabla_{\lambda_{k+1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\alpha\beta\rho\sigma}$ (37)
$-2\sum^{m}_{i=0}{Z}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\lambda\rho\sigma}_{(i)}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}^{\nu}_{~\lambda\rho\sigma}+{g}^{\mu\rho}{g}^{\nu\sigma}\frac{\partial{L}}{\partial{g}^{\rho\sigma}}\,.$
By the comparison between Eqs. (31) and (37), we have proved by performing direct computation that there exists the identity expressing the linear combination of the three tensors ${W}^{\mu\nu}$, ${B}^{\mu\nu}$ and ${P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}$ in the following manner
${W}^{\mu\nu}$ $=$ ${B}^{\mu\nu}-2{P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}+2\sum^{m}_{i=0}{Z}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\lambda\rho\sigma}_{(i)}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}^{\nu}_{~\lambda\rho\sigma}$ (38)
$+\frac{1}{2}\sum^{m}_{i=0}\sum^{i}_{k=1}{Z}^{\lambda_{1}\cdot\cdot\cdot\lambda_{k-1}\mu\lambda_{k+1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}_{(i)}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{k-1}}\nabla^{\nu}\nabla_{\lambda_{k+1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\alpha\beta\rho\sigma}\,.$
Besides, it will be shown by Eqs. (78) and (79) below that the identity (38) can be alternatively derived out of Eq. (7) with the substitution $\delta\rightarrow\mathcal{L}_{\zeta}$.
Note that Eq. (37) is not the simplest form that $(\Delta{E})^{\mu\nu}$ actually belongs to. As a matter of fact, one is able to go further. To see this, like in the work [6], it is feasible to resort to the Lie derivative of the Lagrangian density L{L} along the vector field $\zeta^{\mu}$. In practice, this is able to be obtained by means of the setting that the variation of the Lagrangian density L{L} is performed via the Lie derivative $\mathcal{L}_{\zeta}$ acting on L{L}. Accordingly, replacing the $\delta$ operator in the variation equation (2) of the Lagrangian density with $\mathcal{L}_{\zeta}$ gives rise to
$\boxed{\mathcal{L}_{\zeta}{L}=\frac{\partial{L}}{\partial{g}^{\mu\nu}}\mathcal{L}_{\zeta}{g}^{\mu\nu}+\sum^{m}_{i=0}\frac{\partial{L}}{\partial\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}}\,.$ (39)
That is to say, equation (39) can be understood as an outcome arising from the variation of the Lagrangian density, or that for the Lagrangian in the manner $\mathcal{L}_{\zeta}\mathcal{L}=\sqrt{-g}(\mathcal{L}_{\zeta}{L}+{L}\nabla_{\mu}\zeta^{\mu})$. Here we point out that Eq. (39) is sufficient for our purpose. It is unnecessary for us to introduce the other expression for $\mathcal{L}_{\zeta}{L}$ that consists of the covariant derivative of the Lagrangian density L{L} (see Eq. (41) below for its explicit expression) as shown in Ref. [6]. According to the rules of Lie derivative acting on tensors, we observe that equation (39) can be further recast into
$\zeta^{\nu}\left(\nabla_{\nu}{L}-\sum^{m}_{i=0}Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}\nabla_{\nu}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\alpha\beta\rho\sigma}\right)+2(\Delta{E})^{\mu\nu}\nabla_{\mu}\zeta_{\nu}=0\,.$ (40)
The above equation holds for any vector field $\zeta^{\mu}$. As a result of this, there have to exist the following two identities, given respectively by
$\nabla_{\nu}{L}$ $=$ $\sum^{m}_{i=0}Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}\nabla_{\nu}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\alpha\beta\rho\sigma}\,$, (41)
$(\Delta{E})^{\mu\nu}$ $=$ $0\,.$ (42)
Equation (41) stands for the expression for the covariant derivative of the Lagrangian density expanded in the variables $\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\alpha\beta\rho\sigma}$s $(i=0,1,\cdot\cdot\cdot,m)$. Unlike in Ref. [6], we here derive it out of the Lie derivative of the Lagrangian density rather than treat it as the other condition to produce the identity (42) in addition to the one given by Eq. (39). That is to say, within the context of Padmanabhan’s method, equation (39) is actually the sufficient condition for the derivation of the identity (42), which is adequately responsible for the elimination of the derivative of the Lagrangian density with respect to the metric appearing in the field equations. Moreover, as shown in Eq. (74), the identity (41) will play an important role in proving the vanishing of the divergence for the expression of the field equations by carrying out direct computation.
From Eq. (42), we directly obtain the following identity
${g}^{\mu\rho}{g}^{\nu\sigma}\frac{\partial{L}}{\partial{g}^{\rho\sigma}}$ $=$ $\frac{1}{2}\sum^{m}_{i=0}\sum^{i}_{k=1}{Z}^{\lambda_{1}\cdot\cdot\cdot\lambda_{k-1}\mu\lambda_{k+1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}_{(i)}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{k-1}}\nabla^{\nu}\nabla_{\lambda_{k+1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\alpha\beta\rho\sigma}$ (43)
$+2\sum^{m}_{i=0}{Z}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\lambda\rho\sigma}_{(i)}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}^{\nu}_{~\lambda\rho\sigma}$
$=$ $2{P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}+{W}^{\mu\nu}-{B}^{\mu\nu}\,.$
The above identity associates the derivative of the Lagrangian density with respect to the metric with the one with regard to the covariant derivatives of the Riemann tensor. As is what will be demonstrated in Eq. (80) below, the equality (43) can be reproduced in an alternative way. Apart from Eq. (43), owing to the fact that the second-rank tensors ${B}^{\mu\nu}$ and $\frac{\partial{L}}{\partial{g}^{\mu\nu}}$ are symmetric, $(\Delta{E})^{[\mu\nu]}=0$ appearing as a result from Eq. (42) leads to another identity
${W}^{[\mu\nu]}=-2{P}^{[\mu|\lambda\rho\sigma|}{R}^{\nu]}_{~~\lambda\rho\sigma}\,$, (44)
or equivalently,
$\sum^{m}_{i=0}\sum^{i}_{k=1}{Z}^{\lambda_{1}\cdot\cdot\cdot\lambda_{k-1}[\mu|\lambda_{k+1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma|}_{(i)}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{k-1}}\nabla^{\nu]}\nabla_{\lambda_{k+1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\alpha\beta\rho\sigma}$
$+4\sum^{m}_{i=0}{Z}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}[\mu|\lambda\rho\sigma|}_{(i)}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}^{\nu]}_{~~\lambda\rho\sigma}=0\,.$ (45)
Equation (44) or (45) reveals that the second-rank tensor ${X}^{\mu\nu}$ is symmetric, resulting in $\tilde{E}^{\mu\nu}=\tilde{E}^{\nu\mu}$. In particular, when the Lagrangian (1) depends on ${g}^{\mu\nu}$ and $\Box^{i}{R}_{\mu\nu\rho\sigma}$s, both the two identities (43) and (45) coincide with the corresponding results given by [23]. What is more, as a direct consequence of Eq. (42), we have proved strictly that
${X}_{\mu\nu}=\frac{\delta{L}}{\delta{g}^{\mu\nu}}\,.$ (46)
Equation (46) yields in a straightforward way that both ${E}^{\mu\nu}$ and $\tilde{E}^{\mu\nu}$ are completely identified with each other, that is,
${E}^{\mu\nu}=\tilde{E}^{\mu\nu}={P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}-2\nabla_{\rho}\nabla_{\sigma}{P}^{\rho\mu\nu\sigma}-\frac{1}{2}{L}g^{\mu\nu}+{W}^{\mu\nu}\,.$ (47)
This implies that we verify once again the consequence ${E}^{\mu\nu}=\tilde{E}^{\mu\nu}$ obtained in our previous work [20] in a different way. From equation (47), it is indicated that it is unnecessary to distinguish ${E}^{\mu\nu}$ and $\tilde{E}^{\mu\nu}$ in essence, while the tilde over the expression for the field equations only represents that this expression is derived out of the surface term. Nonetheless, one observes that employing ${X}_{\mu\nu}$ as a replacement for $\frac{\delta{L}}{\delta{g}^{\mu\nu}}$ completely avoids the calculation on Eq. (43) and this makes $\tilde{E}^{\mu\nu}$ naturally eliminate the tensor $\frac{\partial{L}}{\partial{g}^{\mu\nu}}$ so that $\tilde{E}^{\mu\nu}$ takes a simpler form than ${E}^{\mu\nu}$ given by Eq. (16). Besides, the composition manner of $\tilde{E}^{\mu\nu}$ provides convenience for verifying the vanishing for the covariant divergence of the expression of the field equations via straightforward computations as is what will be demonstrated in Section 6. Moreover, some properties on the trace of $\tilde{E}^{\mu\nu}$ can be found in [25].
In the above, it has been proved that the expression $\tilde{E}^{\mu\nu}$ for the field equations appears as a consequence from the surface term. However, due to the fact that equation (43) allows for the removal of the tensor $\frac{\partial{L}}{\partial{g}^{\mu\nu}}$ or $\frac{\partial{L}}{\partial{g}_{\mu\nu}}$ in the field equations, $\tilde{E}_{\mu\nu}$ is also able to be thought of as the result arising from the substitution of Eq. (43) into the Euler-Lagrange expression ${E}_{\mu\nu}$ given by Eq. (16), that is,
$\tilde{E}_{\mu\nu}={E}_{\mu\nu}\left(\frac{\partial{L}}{\partial{g}^{\mu\nu}}\rightarrow 2{P}_{\mu}^{~\lambda\rho\sigma}{R}_{\nu\lambda\rho\sigma}+{W}_{\mu\nu}-{B}_{\mu\nu}\right)\,.$ (48)
In fact, this is fully in accordance with the spirit of the procedure proposed by Padmanabhan to derive the equations of motion associated to the theories of gravity with the Lagrangian only built out of the metric and its Riemann tensor within Ref. [6] (the key point of Padmanabhan’s method lies within Eq. (61) or the first equivalence relation of Eq. (63) in Section 5). To this point, $\tilde{E}^{\mu\nu}$ can be regarded as the most natural generalization for the expression of the equations of motion given by Ref. [6] to the theories of gravity involving the arbitrary order covariant derivatives of the Riemann tensor. Consequently, just from the perspective for the derivation of the equations of motion, our procedure for the extraction of the field equations out of the surface term provides another path towards the reproduction and the generalization for the field equations via Padmanabhan’s method, while the former offers an additional advantage of ruling out the calculations involving the derivative of the Lagrangian density with respect to the metric tensor.
Finally, apart from the equations of motion for the gravitational field, we demonstrate that the surface term (25) also encodes the information for conserved current. To see this, we turn our attention to extract the conserved current from such a term. Substituting $X^{\mu\nu}={E}^{\mu\nu}+\frac{1}{2}{g}^{\mu\nu}L$ into Eq. (25), we obtain the conserved current ${J}^{\mu}$ associated to the vector field $\zeta^{\mu}$, defined in terms of the quantity $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ through
${J}^{\mu}=2\zeta_{\nu}{E}^{\mu\nu}+\zeta^{\mu}{L}-\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})\,$, (49)
or equivalently defined in terms of the anti-symmetric tensor ${Q}^{\mu\nu}$ as ${J}^{\mu}=\nabla_{\nu}{Q}^{\mu\nu}$. Actually, here the current ${J}^{\mu}$ is nothing else but the conventional off-shell Noether current relative to the vector field $\zeta^{\mu}$ [18, 26, 27], while the anti-symmetric rank-two tensor ${Q}^{\mu\nu}$ is the corresponding Noether charge 2-form [1, 22]. We here present the explicit form for the Noether charge 2-form within the context of the diffeomorphism invariant theories of gravity relying on the covariant derivatives of the Riemann tensor. This could be helpful to understand the Clausius relation relative to the diffeomorphism invariant gravities encompassing higher derivative and higher curvature terms [17, 18].
4 The comparison between two ways of identifying the expressions for the field equations
We have illustrated in great detail that the expression for the field equations can be derived out of the surface term arising from the variation of the Lagrangian in the previous section, as well as in our previous work [20]. The main process to do so in both the works is the same, which consists of three steps:
[…]
The only difference appearing in the last step is the manner in which we verify that both the expressions $\tilde{E}^{\mu\nu}$ and ${E}^{\mu\nu}$ for the field equations indeed coincide with each other. Owing to this, within the present section, we will offer a detailed comparison between the methods to identify $\tilde{E}^{\mu\nu}$ with ${E}^{\mu\nu}$ employed respectively in the present work and Ref. [20]. Meanwhile, motivated by the results, we will reveal several connections between the expression for the field equations and the off-shell Noether current.
In the previous section, it has been demonstrated that the relationship $\tilde{E}^{\mu\nu}={E}^{\mu\nu}$ directly results from Eq. (46), which arises as a consequence of the pivotal relationship among the symmetric rank-two tensor $\frac{\partial{L}}{\partial{g}^{\mu\nu}}$ and the derivative of the Lagrangian density with respect to the $i$th-order covariant derivatives of the Riemann tensor $\frac{\partial{L}}{\partial\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}}$s, given by Eq. (43). This equation, resulting from Eq. (39), actually serves as the starting point of identifying $\tilde{E}^{\mu\nu}$ with ${E}^{\mu\nu}$, and it provides a solid foundation for establishing the connection $\tilde{E}^{\mu\nu}={E}^{\mu\nu}$ since it arises as a straightforward outcome for the Lie derivative of the Lagrangian density along an arbitrary smooth vector field without any other conditions. At last, the off-shell Noether current ${J}^{\mu}$ given by Eq. (49) appears as a result of $\tilde{E}^{\mu\nu}={E}^{\mu\nu}$. To sum up, the aforementioned process from the perspective of computing directly the difference between ${E}^{\mu\nu}$ and $\tilde{E}^{\mu\nu}$ goes as follows:
[…]
Apart from the above process for the handling of $\tilde{E}^{\mu\nu}={E}^{\mu\nu}$, it has been illustrated in our previous work [20] that such an equality can be obtained from the off-shell Noether current perspective. To do so, conversely, the off-shell Noether current ${J}^{\mu}$ is adopted to produce the expression for the field equations rather than arises as the outcome of this expression. As is shown by Eq. (103), the covariant divergence of this current is associated to the identity
$\boxed{\nabla_{\mu}{J}^{\mu}=2\zeta_{\nu}\nabla_{\mu}{E}^{\mu\nu}}\,$, (52)
directly originating from the variation equation (15) of the Lagrangian under the transformation $\delta\rightarrow\mathcal{L}_{\zeta}$. As a matter of fact, the identity (52), which can be viewed as a bridge between the divergence for the conserved current and that for the expression of the field equations, constitutes the starting point to yield $\tilde{E}^{\mu\nu}={E}^{\mu\nu}$ in the work [20]. By contrast, employing the identity (52) instead of Eq. (43) as the starting point is relatively simple. Besides, in that work, equation (43) is treated as the consequence resulting from $\tilde{E}^{\mu\nu}={E}^{\mu\nu}$ rather than the foundation to verify this equality like in the present work. In short, the process to obtain $\tilde{E}^{\mu\nu}={E}^{\mu\nu}$ from the off-shell Noether current perspective can be summarized as
[…]
which can be regarded as the inverse of Eq. (51).
Actually, according to Eq. (109), the form (25) for the surface term $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ results in $\nabla_{\mu}{E}^{\mu\nu}=0$ under the condition (52). That is to say, equation (52) can be replaced with the conserved equation $\nabla_{\mu}{J}^{\mu}=0$ for the Noether current ${J}^{\mu}$. Therefore, for the Lagrangian (1) accommodating diffeomorphism invariance, on the basis of Eqs. (51) and (53), there exist the following equivalence relations:
[…]
From Eq. (54), we observe that both the two ways to verify ${E}^{\mu\nu}=\tilde{E}^{\mu\nu}$ presented respectively by this work and our previous one [20] are equivalent.
Furthermore, we concentrate on the relations between the conserved Noether current and the expression for the field equations. With the help of the correspondence
$\nabla_{\mu}{J}^{\mu}=0\quad\Leftrightarrow\quad\nabla_{\mu}{E}^{\mu\nu}=0\,$ (55)
arising from the identity (52), in terms of Eq. (54), one further builds the equivalence relationships among the generalized Bianchi identity $\nabla_{\mu}{E}^{\mu\nu}=0$, the conserved equation $\nabla_{\mu}{J}^{\mu}=0$, the expression for the field equations and the off-shell Noether current, given by
$\boxed{\nabla_{\mu}{J}^{\mu}=0\quad\Leftrightarrow\quad\nabla_{\mu}{E}^{\mu\nu}=0\quad\Leftrightarrow\quad{E}^{\mu\nu}={X}^{\mu\nu}-\frac{1}{2}{L}{g}^{\mu\nu}\quad\Leftrightarrow\quad{J}^{\mu}=\nabla_{\nu}{Q}^{\mu\nu}}\,.$ (56)
All the equivalence relationships appearing in Eq. (56) are displayed by Fig. 1. That is to say, the expression for the equations of motion can arise as the result of the conserved equation for the current. Apart from this, the equivalence between $\nabla_{\mu}{E}^{\mu\nu}=0$ and ${E}^{\mu\nu}={X}^{\mu\nu}-\frac{1}{2}{L}{g}^{\mu\nu}$ in Eq. (56) allows for the derivation for the expression of the field equations out of the vanishing of its divergence. In light of this, the equivalence relationships given by Eq. (56) might provide some insights on the thermodynamic derivation for the gravitational field equations [17, 18, 28]. Particularly, we propose that the expression for the field equations stemming from the thermodynamic derivation should be $\tilde{E}^{\mu\nu}$ ranther than ${E}^{\mu\nu}$. It is worth mentioning that all the three equations (25), (39) and (52) play a crucial role in establishing the equivalence relationships appearing in Eqs. (54) and (56). In fact, as is what will be illustrated in detail within the next section, since equation (25) holds identically, all the three equations constantly coexist as long as either of the two equations (39) and (52) among them comes into existence.
Figure 1: The equivalence relationships associated to the expression ${E}^{\mu\nu}$ for the field equations and the conserved current ${J}^{\mu}$ are unveiled. The definition (49) for the conserved current is employed. The expression (25) for the surface term is instrumental in establishing these equivalence relationships. Additionally, equation (52), which is equivalent to equation (39), gets involved in establishing the equivalence between $\nabla_{\mu}{E}^{\mu\nu}=0$ and $\nabla_{\mu}{J}^{\mu}=0$.
At the end, we observe that equation (56) gives rise to another interesting relation between the conserved current and the expression for the field equations, that is, the structure for the equivalence relationship associated to the off-shell Noether current
$\nabla_{\mu}{J}^{\mu}=0\quad\Leftrightarrow\quad{J}^{\mu}=\nabla_{\nu}{Q}^{\mu\nu}\,$ (57)
bears a resemblance to the one involving the expression for the field equations
$\nabla_{\mu}{E}^{\mu\nu}=0\quad\Leftrightarrow\quad{E}^{\mu\nu}={X}^{\mu\nu}-\frac{1}{2}{L}{g}^{\mu\nu}\,.$ (58)
Both the equations (57) and (58) establish the equivalence between the concrete expression of a quantity and the vanishing of its covariant divergence. In light of Eqs. (55), (57) and (58), together with
${E}^{\mu\nu}={X}^{\mu\nu}-\frac{1}{2}{L}{g}^{\mu\nu}\quad\Leftrightarrow\quad{J}^{\mu}=\nabla_{\nu}{Q}^{\mu\nu}\,$, (59)
we propose that there exists the ${E}^{\mu\nu}/{J}^{\mu}$ correspondence between the expression ${E}^{\mu\nu}$ for the equations of motion and the conserved current ${J}^{\mu}$ in the framework of diffeomorphism invariant theories of pure gravity.
5 Relations among the three equations (25), (39) and (52)
Due to the analysis performed in the previous two sections, the combination of the two equations (25) and (39) plays a critical role in the derivation for the field equations without the derivative of the Lagrangian density with respect to the metric in the present work, while it was illustrated in our previous work [20] that the expression for the equations of motion can also appear as a consequence from the combination of Eqs. (25) and (52), that is,
[…]
Apparently, the three equations (25), (39) and (52) that arise from the substitution of the variation operator by the Lie derivative along an arbitrary smooth vector field occupy an important position in our study. Thus, within this section, we will reveal some interesting and significant relations among them. Inspired with the results, we will discuss the main implications of Padmanabhan’s method.
Our fundamental purpose for resorting to the Lie derivative of the Lagrangian density L{L} along the diffeomorphism generator $\zeta^{\mu}$ given by Eq. (39) is to establish the connection between $\frac{\partial{L}}{\partial{g}^{\mu\nu}}$ and $\frac{\partial{L}}{\partial\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}}$s, by means of which the rank-two tensor $\frac{\partial{L}}{\partial{g}^{\mu\nu}}$ in the Euler-Lagrange equations can be eliminated. It has been demonstrated that Eq. (39) gives rise to Eq. (43) that indeed assists us to achieve this goal. In fact, equation (39) is completely equivalent to the combination of both the equations (41) and (43). Moreover, by the aid of Eq. (74), equation (41) is able to arise as a consequence from the generalized Bianchi identity $\nabla_{\mu}{E}^{\mu\nu}=0$ written as $\nabla^{\nu}{L}=2\nabla_{\mu}{X}^{\mu\nu}$ and vice versa, while Eq. (54) shows that Eq. (43) is equivalent to ${E}^{\mu\nu}=\tilde{E}^{\mu\nu}$. Consequently, we obtain
[…]
As shown by Eq. (61), there exist two ways to verify Eq. (39). One is to confirm that both the two equations (41) and (43) hold through direct calculations, and the other is to test ${E}^{\mu\nu}={X}^{\mu\nu}-\frac{1}{2}{g}^{\mu\nu}{L}$ together with the generalized Bianchi identity for ${E}^{\mu\nu}$.
What is more, according to Eq. (106), under the condition that the surface term $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ possesses the form (25), the identity (52) gives rise to $\nabla_{\mu}{E}^{\mu\nu}=0$ and ${E}^{\mu\nu}=\tilde{E}^{\mu\nu}$ simultaneously. The former leads to Eq. (41), and the latter yields Eq. (43) in terms of Eq. (53). Ultimately, both of Eqs. (41) and (43) jointly produce Eq. (39). On the other hand, by virtue of Eq. (61), it is easy to see that ${E}^{\mu\nu}=\tilde{E}^{\mu\nu}$ together with $\nabla_{\mu}{E}^{\mu\nu}=0$ can be derived out of Eq. (39). Then the combination of ${E}^{\mu\nu}=\tilde{E}^{\mu\nu}$ and $\nabla_{\mu}{E}^{\mu\nu}=0$ further leads to Eq. (52) in the form
$\nabla_{\mu}\big(\zeta^{\mu}{L}\big)$ $=$ $-2{E}^{\mu\nu}\nabla_{\mu}\zeta_{\nu}+\nabla_{\mu}\big(2\zeta_{\nu}{X}^{\mu\nu}\big)$ (62)
$=$ $-2{E}^{\mu\nu}\nabla_{\mu}\zeta_{\nu}+\nabla_{\mu}\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})\,.$
Equation (62) is just Eq. (15) with $\delta$ substituted by $\mathcal{L}_{\zeta}$. In order to arrive at the second equality in the above equation, here we have utilized the relation $\nabla_{\mu}\big(2\zeta_{\nu}{X}^{\mu\nu}\big)=\nabla_{\mu}\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ resulting from Eq. (25). Hence, it is illustrated that Eq. (52) can be reproduced by Eq. (39) with the help of Eq. (25). Due to the aforementioned analysis, as long as $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ is provided with the form (25) or its covariant divergence coincides with that for the vector $2\zeta_{\nu}{X}^{\mu\nu}$, we have
[…]
It is worth pointing out that the first equivalence relationship in Eq. (63) holds without the requirement to impose the condition that $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ admits the form (25), which only plays a role in the establishment of the second equivalence relationship. As a matter of fact, equation (61) or the first equivalence relationship in Eq. (63) harbours the main implications of the method proposed by Padmanabhan for the derivation of the field equations in the absence of the derivative of the Lagrangian density with respect to the metric tensor [6]. To this point, equation (61) could be understood as the foundation of this method. Apparently, in terms of Eq. (63), we have demonstrated that each of the two equations (39) and (52) can be chosen as the sufficient condition for identifying $\tilde{E}^{\mu\nu}$ with ${E}^{\mu\nu}$, and the spirit for the procedure using the former to yield ${E}^{\mu\nu}=\tilde{E}^{\mu\nu}$ is in accordance with that for Padmanabhan’s method. Nevertheless, employing the latter to verify ${E}^{\mu\nu}=\tilde{E}^{\mu\nu}$ possesses the advantage of getting rid of the derivative of the Lagrangian density with respect to the metric during the whole process.
On the basis of Eq. (63), let us delve deeper into the association of Eq. (25) with both the equations (39) and (52). According to the first equivalence relationship in Eq. (63) together with Eq. (119), the combination of the generalized Bianchi identity $\nabla_{\mu}{E}^{\mu\nu}=0$ resulting from Eq. (39) with Eq. (52) enables one to arrive at
[…]
Here the rank-two anti-symmetric tensor contained in the covariant divergence term obeys $\nabla_{\nu}(\text{{anti-symmetric tensor}})^{\mu\nu}=\nabla_{\nu}{Q}^{\mu\nu}$. Apparently, according to Eq. (64), for the surface term $\Theta^{\mu}$ arising from the variation of the Lagrangian, the validity for the two equations (39) and (52) constitutes the sufficient condition for expressing $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ as the form (25). However, this can be simplified further. Specifically, note that both the two equations (39) and (52) originate from the replacement of the variation operator $\delta$ with the Lie derivative $\mathcal{L}_{\zeta}$ within the variation equation of the Lagrangian (1) just expressed respectively as two equivalent but formally different equations (5) and (15). To this point, it is reasonable to conclude that the equivalence between Eqs. (5) and (15) results in the coexistence of Eqs. (39) and (52). As a matter of fact, such a conclusion is verified to hold indeed by means of direct calculations within Appendix B. Therefore, from the perspective for the coexistence of Eqs. (39) and (52), the validity for either of them can be fully responsible for the decomposition of the surface term $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ into the form (25). On the other hand, regardless of the aforementioned derivation for Eq. (25), in terms of our process to derive this equation given by Section 3, it has actually been shown that Eq. (25) holds identically within the context of the Lagrangian (1). Moreover, the existence of this equation can be independent of Eqs. (39) and (52), attributed to the fact that Eq. (25) directly arises from the substitution of $\delta$ with $\mathcal{L}_{\zeta}$ in the surface term $\Theta^{\mu}$, which merely comes from the sum of the scalar $Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ with respect to $i$ from 0 to $m$ appearing in Eq. (2).
What is more, due to the analysis mentioned above on the existence for the form (25), it is really unnecessary to particularly stress on that the structure of $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ plays a role in establishing the equivalence relation between Eqs. (39) and (52) as before. Apparently, as a consequence of Eqs. (63) and (64), the setting up of the two equations (39) and (52) automatically brings about the inference that they are completely equivalent to each other (i.e. existence implies equivalence), namely,
[…]
From Eq. (65), by utilizing Eqs. (63) and (64), we actually have presented a strict proof that the equivalence relationship between the two forms (5) and (15) for the variation of the Lagrangian (1) is still reserved after the variation operator in them is substituted with the Lie derivative along an arbitrary vector field. Another proof for Eq. (65) in terms of straightforward calculations can be found within Appendix B. The equivalence between Eqs. (39) and (52) causes that the equations of motion arising from Padmanabhan’s method have to be in agreement with our ones from the surface term perspective.
Finally, on the basis of the aforementioned analysis in the present section, we turn to summarizing the relationships among the three equations (25), (39) and (52). They are established under the condition that the variation operator in Eqs. (18), (5) and (15) is allowed to be replaced with the Lie derivative $\mathcal{L}_{\zeta}$, respectively. In particular, as has been shown in Appendix B, both the two equations (39) and (52) always coexist attributed to the equivalence between Eqs. (5) and (15), and they are equivalent under their coexistence. Besides, the combination of Eqs. (25) and (52) is completely equivalent to the conserved equation for the Noether current (49) plus the generalized Bianchi identity for the Euler-Lagrange expression (16), that is,
[…]
Obviously, here Eq. (66) provides another way to confirm the validity of Eqs. (25) and (52). Moreover, there exists the fact that allowing the substitution $\delta\rightarrow\mathcal{L}_{\zeta}$ in any two of Eqs. (5), (15) and (18) guarantees that all the three equations (25), (39) and (52) must come into existence simultaneously. To see this more clearly, on the basis of both the two equations (63) and (64) or according to the two equations (141) and (142), it is easy to find that the three equations (25), (39) and (52) are actually not completely independent to each other and the arbitrary two equations among them can jointly serve as the sufficient condition for yielding the remaining one (see Fig. 2 for illustration). In other words, for the sake of ensuring that all of the three equations are valid, it is only necessary to confirm that the arbitrary two equations of them hold.
Figure 2: Any one of the three equations (25), (39) and (52) is able to result from the rest two. Equation (25) holds identically although it can appear as a consequence from the coexistence of Eqs. (39) and (52). Besides, the fact that equation (25) always holds true offers support for the equivalence between Eqs. (39) and (52).
For instance, in order to guarantee that all the three equations (25), (39) and (52) hold together, it is only required to prove that $\nabla_{\mu}{E}^{\mu\nu}=0$ and $\nabla_{\mu}{J}^{\mu}=0$ by using Eq. (66). Apart from this, only from the perspective of the coexistence relationship between the two equations (39) and (52), a simpler way is to verify that Eq. (39) or (52) is valid since the validity for either of them indicates that the other one is valid as well. In view of Eq. (61), one feasible way to accomplish this is to test directly that both the two equations (41) and (43) hold simultaneously. What is more, it has been demonstrated that the surface term $\Theta^{\mu}$ with $\delta\rightarrow\mathcal{L}_{\zeta}$ can be put into the form (25) via the straightforward calculations having nothing to do with Eqs. (39) and (52). This implies that Eq. (25) can exist independently of these two equations. However, the coexistence of Eqs. (39) and (52) is able to produce Eq. (25). So the existence of Eq. (25) is a necessary condition for guaranteeing that both the equations (39) and (52) hold simultaneously. On the other hand, although the existence for Eq. (25) is not a sufficient condition for determining that either of Eqs. (39) and (52) definitely holds, equation (25) can be employed to assist us to derive either of them out of the other one.
6 The disappearance for the covariant divergence of the expression of the field equations
As is by now well-known, within the framework of general relativity, in order to ensure the conservation law of energy-momentum tensor to hold, the vanishing for the covariant divergence of Einstein tensor is viewed as one of the fundamental assumptions in curved spacetimes. In the same way, the problem for conservation law is of great importance as well within the context of a wide variety of theories for gravity admitting diffeomorphism invariance beyond general relativity. Owing to this, it is necessary for us to take into account the covariant divergence of the expression of the field equations relative to the Lagrangian (1). However, unlike the situation for general relativity, the related calculations become more complicated as a consequence for the appearance of the covariant derivatives of the Rienmann curvature tensor.
Due to the above, it can be found that there exist at least four ways towards the proof for the generalized Bianchi identity $\nabla_{\mu}{E}^{\mu\nu}=0$. (I) According to Eq. (51), the first way is to adopt the identity (43) as the starting point to yield the off-shell Noether current ${J}^{\mu}=\nabla_{\nu}{Q}^{\mu\nu}$, which further gives rise to $\nabla_{\mu}{J}^{\mu}=0$. Then Eq. (52) leads to $\nabla_{\mu}{E}^{\mu\nu}=0$. (II) In light of Eq. (55), the second way is to directly verify the vanishing of the covariant divergence of ${J}^{\mu}$. (III) Under the condition that Eq. (52) holds true, the third way is to utilize the conclusion that the surface term $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ can be expressed as the form (101), directly yielding $\nabla_{\mu}{E}^{\mu\nu}=0$ by means of Eq. (109). (IV) The fourth way is to employ Eq. (39) as the starting point. Such an equation results in two identities (41) and (43). According to Eq. (51), the latter enables one to obtain ${E}^{\mu\nu}=\tilde{E}^{\mu\nu}$. In terms of this, we directly compute the covariant divergence of $\tilde{E}^{\mu\nu}$ to produce $\nabla_{\mu}\tilde{E}^{\mu\nu}=0$ by the aid of the identity (41). Actually, all the four ways are not completely independent to each other, and either of Eqs. (39) and (52) is able to serve as the starting point.
In spite of the fact that the generalized Bianchi identity can be successfully proved via each of the first three ways, in the remainder of this section, we will follow the fourth way to perform straightforward calculations on the covariant divergence of the expression for equations of motion to confirm that it indeed vanishes identically within the context of matter free gravity theories endowed with the Lagrangian (1). All the related calculations are based on the expression (29) for equations of motion arising from the surface term rather than the one (16) corresponding to the Euler-Lagrange equations.
As a beginning, we handle the covariant divergence of the ${X}^{\mu\nu}$ tensor. The straightforward calculations bring forth
$\nabla_{\mu}{X}^{\mu\nu}$ $=$ $R^{\nu}_{~\lambda\rho\sigma}\nabla_{\mu}{P}^{\mu\lambda\rho\sigma}+{P}^{\mu\lambda\rho\sigma}\nabla_{\mu}R^{\nu}_{~\lambda\rho\sigma}-2\nabla_{\mu}\nabla_{\rho}\nabla_{\sigma}{P}^{\rho\mu\nu\sigma}+\nabla_{\mu}{W}^{\mu\nu}\,.$ (67)
Within the above equation, the vector ${P}^{\mu\lambda\rho\sigma}\nabla_{\mu}R^{\nu}_{~\lambda\rho\sigma}$ is alternatively expressed as
${P}^{\mu\lambda\rho\sigma}\nabla_{\mu}R^{\nu}_{~\lambda\rho\sigma}=\frac{1}{2}{P}^{\alpha\beta\rho\sigma}\nabla^{\nu}R_{\alpha\beta\rho\sigma}\,$, (68)
and the vector $\nabla_{\mu}\nabla_{\rho}\nabla_{\sigma}{P}^{\rho\mu\nu\sigma}$ can be rewritten as
$\nabla_{\mu}\nabla_{\rho}\nabla_{\sigma}{P}^{\rho\mu\nu\sigma}=-\nabla_{[\rho}\nabla_{\mu]}\nabla_{\sigma}{P}^{\rho\mu\nu\sigma}=\frac{1}{2}R^{\nu}_{~\lambda\rho\sigma}\nabla_{\mu}{P}^{\mu\lambda\rho\sigma}\,.$ (69)
In order to figure out the divergence of the tensor ${W}^{\mu\nu}=\sum^{m}_{i=0}{W}^{\mu\nu}_{(i)}$, a primary task could be to deal with $\nabla_{\mu}{W}^{\mu\nu}_{(i)}$, yielding
$\nabla_{\mu}{W}^{\mu\nu}_{(i)}$ $=$ $\frac{1}{2}\sum_{k=1}^{i}(-1)^{k-1}\left(\nabla_{\lambda_{k-1}}\cdot\cdot\cdot\nabla_{\lambda_{1}}Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}\right)\nabla^{\nu}\nabla_{\lambda_{k}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\alpha\beta\rho\sigma}$ (70)
$-\frac{1}{2}\sum_{k=1}^{i}(-1)^{k}\left(\nabla_{\lambda_{k}}\cdot\cdot\cdot\nabla_{\lambda_{1}}Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}\right)\nabla^{\nu}\nabla_{\lambda_{k+1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\alpha\beta\rho\sigma}$
[…]
We observe that the vector $R^{\nu}_{~\lambda\rho\sigma}{U}^{\rho\lambda\sigma}_{(i)}$ in Eq. (70) is able to be further expressed as
$R^{\nu}_{~\lambda\rho\sigma}{U}^{\rho\lambda\sigma}_{(i)}$ $=$ $\frac{1}{2}\big(3R^{\nu}_{~[\sigma\rho\lambda]}+2R^{\nu}_{~\lambda\rho\sigma}\big){U}^{\rho\lambda\sigma}_{(i)}$ (71)
$=$ $\frac{1}{2}\big(R^{\nu}_{~\sigma\rho\lambda}+R^{\nu}_{~\lambda\rho\sigma}+R^{\nu}_{~\rho\lambda\sigma}\big){U}^{\rho\lambda\sigma}_{(i)}$
[…]
Then substituting Eq. (71) into (70) leads to
$\nabla_{\mu}{W}^{\mu\nu}_{(i)}=\frac{1}{2}{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}\nabla^{\nu}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\alpha\beta\rho\sigma}-\frac{1}{2}(-1)^{i}\Big(\nabla_{\lambda_{i}}\cdot\cdot\cdot\nabla_{\lambda_{1}}Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}\Big)\nabla^{\nu}R_{\alpha\beta\rho\sigma}\,.$ (72)
As a result of the sum for equation (72) with respect to $i$ from 0 to $m$, the covariant divergence of the ${W}^{\mu\nu}$ tensor is read off as
$\nabla_{\mu}{W}^{\mu\nu}=\frac{1}{2}\sum^{m}_{i=0}{Z}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}_{(i)}\nabla^{\nu}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\alpha\beta\rho\sigma}-\frac{1}{2}P^{\alpha\beta\rho\sigma}\nabla^{\nu}{R}_{\alpha\beta\rho\sigma}\,.$ (73)
In terms of Eqs. (68), (69) and (73), plugging them back into equation (67) gives rise to
$\nabla_{\mu}{X}^{\mu\nu}$ $=$ $\frac{1}{2}\sum^{m}_{i=0}{Z}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}_{(i)}\nabla^{\nu}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\alpha\beta\rho\sigma}\,.$ (74)
Here we point out that both the two equations (73) and (74) hold identically and they depend upon the ${Z}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}_{(i)}$ tensors.
With Eq. (74) at hand, combining it with equation (41) assists us to arrive ultimately at the Bianchi-type identity
$\nabla_{\mu}{E}^{\mu\nu}=\nabla_{\mu}\tilde{E}^{\mu\nu}=\nabla_{\mu}{X}^{\mu\nu}-\frac{1}{2}\nabla^{\nu}{L}=0\,.$ (75)
Consequently, we prove that ${E}^{\mu\nu}$ is indeed divergence-free.
Finally, we point out that the vector $\nabla_{\mu}{W}^{\mu\nu}$ can be derived out of Eq. (7) under the replacement $\delta\rightarrow\mathcal{L}_{\zeta}$, that is,
$Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ $=$ $(-1)^{i}\big(\nabla_{\lambda_{i}}\cdot\cdot\cdot\nabla_{\lambda_{1}}Z_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\big)\mathcal{L}_{\zeta}{R}_{\mu\nu\rho\sigma}$ (76)
$-2{B}^{\mu\nu}_{(i)}\nabla_{\mu}\zeta_{\nu}+\nabla_{\mu}\bar{\Theta}^{\mu}_{(i)}(\delta\rightarrow\mathcal{L}_{\zeta})\,.$
Specifically, substituting Eq. (22) into (76) and summing the result with respect to $i$, we have
$2\zeta_{\nu}\nabla_{\mu}{W}^{\mu\nu}$ $=$ $2\big({B}^{\mu\nu}-{W}^{\mu\nu}\big)\nabla_{\mu}\zeta_{\nu}-P^{\alpha\beta\rho\sigma}\mathcal{L}_{\zeta}{R}_{\alpha\beta\rho\sigma}$ (77)
$+\sum^{m}_{i=0}{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}\,.$
The above equation can be recast into
$\zeta_{\nu}\left(2\nabla_{\mu}{W}^{\mu\nu}+{P}^{\alpha\beta\rho\sigma}\nabla^{\nu}{R}_{\alpha\beta\rho\sigma}-\sum^{m}_{i=0}{Z}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}_{(i)}\nabla^{\nu}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\alpha\beta\rho\sigma}\right)+\Upsilon^{\mu\nu}\nabla_{\mu}\zeta_{\nu}=0\,$, (78)
where the rank-two tensor $\Upsilon^{\mu\nu}$ is defined through
$\Upsilon^{\mu\nu}$ $=$ $-\sum^{m}_{i=0}\sum^{i}_{k=1}{Z}^{\lambda_{1}\cdot\cdot\cdot\lambda_{k-1}\mu\lambda_{k+1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}_{(i)}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{k-1}}\nabla^{\nu}\nabla_{\lambda_{k+1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\alpha\beta\rho\sigma}$ (79)
$-4\sum^{m}_{i=0}{Z}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\lambda\rho\sigma}_{(i)}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}^{\nu}_{~\lambda\rho\sigma}+4{P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}+2{W}^{\mu\nu}-2{B}^{\mu\nu}\,.$
Within Eq. (78), since this equation is valid for any vector field $\zeta^{\mu}$, the vector field proportional to $\zeta^{\mu}$ and the tensor $\Upsilon^{\mu\nu}$ have to vanish together. The former conclusion gives rise to Eq. (73) for $\nabla_{\mu}{W}^{\mu\nu}$, while $\Upsilon^{\mu\nu}=0$ reproduces the equality (38). On the other hand, by means of the combination of Eq. (39) with (77), we obtain
$\zeta_{\nu}\left(2\nabla_{\mu}{W}^{\mu\nu}+{P}^{\alpha\beta\rho\sigma}\nabla^{\nu}{R}_{\alpha\beta\rho\sigma}-\nabla^{\nu}{L}\right)$
$+2\left(2{P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}+{W}^{\mu\nu}-{B}^{\mu\nu}-\frac{\partial{L}}{\partial{g}^{\rho\sigma}}{g}^{\mu\rho}{g}^{\nu\sigma}\right)\nabla_{\mu}\zeta_{\nu}=0\,.$ (80)
In view of the arbitrariness of the vector field $\zeta^{\mu}$ in Eq. (80), another form for $\nabla_{\mu}{W}^{\mu\nu}$, resulting from the vanishing of the vector contracted with $\zeta_{\nu}$, is read off as
$\nabla_{\mu}{W}^{\mu\nu}=\frac{1}{2}\nabla^{\nu}{L}-\frac{1}{2}{P}^{\alpha\beta\rho\sigma}\nabla^{\nu}{R}_{\alpha\beta\rho\sigma}\,.$ (81)
Meanwhile, apart from Eq. (81), the disappearance of the rank-two tensor contracted with $\nabla_{\mu}\zeta_{\nu}$ leads to the equality (43) once again.
7 The procedure for extracting the expression of the field equations out of the surface term
According to the aforementioned analysis, equation (63) enables us to conclude that the validity for either of Eqs. (39) and (52) is a sufficient and necessary condition for that (I) the ${X}^{\mu\nu}$ tensor arising from the surface term can be employed to define the expression ${E}^{\mu\nu}$ for the equations of motion as ${E}^{\mu\nu}={X}^{\mu\nu}-\frac{1}{2}{L}{g}^{\mu\nu}$ and (II) the covariant divergence of the resulting ${E}^{\mu\nu}$ satisfies $\nabla_{\mu}{E}^{\mu\nu}=0$.
In terms of the conclusion mentioned above together with Eq. (54), the process (50) devoted to the derivation for the expression of the field equations from the surface term perspective is concretized into the following.
Step 1: Varying the Lagrangian density L{L} instead of the Lagrangian $\mathcal{L}$ with respect to all the variables gives rise to Eq. (2). Note that such an equation describing the variation for the Lagrangian density constitutes the starting point of the entire process. Subsequently, on the basis of Eq. (2), regardless of any term associated with the variation of the metric tensor, handling all the terms involving the variation of the variables $\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$s $(i=0,\cdot\cdot\cdot,m)$ leads to the surface term $\Theta^{\mu}$.
Step 2: In terms of the surface term $\Theta^{\mu}$, under the transformation that the variation operator $\delta$ is replaced with the Lie derivative $\mathcal{L}_{\zeta}$ along an arbitrary vector field $\zeta^{\mu}$, this surface term is able to be decomposed into the form (25), which generates the rank-two tensor ${X}^{\mu\nu}$ together with the anti-symmetric one ${Q}^{\mu\nu}$. As a matter of fact, decomposing the surface term with $\delta\rightarrow\mathcal{L}_{\zeta}$ into the form given by Eq. (101) can be regarded as the crucial step towards extracting the expression for the equations of motion.
Step 3: In terms of the expression (25) for the surface term $\Theta^{\mu}$ with $\delta\rightarrow\mathcal{L}_{\zeta}$, under the condition that equation (2) with $\delta$ substituted by $\mathcal{L}_{\zeta}$, presented by Eq. (39), is supposed to be valid, or under the condition that equation (52) holds, the expression ${E}^{\mu\nu}$ for the field equations derived out of the variation of the Lagrangian can be built from the ${X}^{\mu\nu}$ tensor, written as ${E}^{\mu\nu}={X}^{\mu\nu}-\frac{1}{2}{L}{g}^{\mu\nu}$. Here the validity for either of Eqs. (39) and (52) supplies a guarantee of the symmetry for the rank-two tensor ${X}^{\mu\nu}$, namely, ${X}^{\mu\nu}={X}^{\nu\mu}$. Meanwhile, the combination of ${E}^{\mu\nu}$ with the quantity $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ leads to the Noether current ${J}^{\mu}$, taking the form ${J}^{\mu}=2\zeta_{\nu}{E}^{\mu\nu}+\zeta^{\mu}{L}-\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ or ${J}^{\mu}=\nabla_{\nu}{Q}^{\mu\nu}$.
Moreover, inspired with the three steps mentioned above, we find that the variation of the Lagrangian density together with its Lie derivative along an arbitrary vector field can be sufficient for producing the expressions of the field equations and the Noether current in a uniform framework. In terms of this, as an application to fully illustrate the surface term perspective towards the derivation for the expression of the field equations proposed above, we demonstrate within the remainder of this section that all the results related to the Lagrangian (1) can be directly generalized to the following Lagrangian
$\sqrt{-g}{L}_{gen}=\sqrt{-g}{L}_{gen}\left({g}^{\mu\nu},{R}_{\alpha\beta\rho\sigma},\nabla_{\lambda}\right)\,.$ (82)
Here the Lagrangian (82) is more general than the one given by Eq. (1) in form. However, each of them can be employed to describe the theories of gravity depending on the metric and the covariant derivatives of the Riemann tensor.
Specifically, for such gravity theories armed with the Lagrangian (82), as a starting point, we only need to compute the variation of the Lagrangian density ${L}_{gen}$. In general, the result can be always put into the following form
$\delta{L}_{gen}=\check{S}_{\mu\nu}\delta{g}^{\mu\nu}+\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}\,.$ (83)
Within Eq. (83), the nonnegative integer $m$ stands for the maximum number of the covariant derivative operator acting on the Riemann curvature tensor, while the symmetric rank-two tensor $\check{S}_{\mu\nu}$ and the rank-$(i+4)$ tensors $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$s are actually allowed to be arbitrary in form although they are constrained by the variation of the Lagrangian density. For convenience, like before, we impose the restriction that the last four indices in the $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$ tensor identically satisfy $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}=\Delta^{\mu\nu\rho\sigma}_{\alpha\beta\gamma\lambda}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\gamma\lambda}$ attributed to the relation $\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}=\Delta^{\alpha\beta\gamma\lambda}_{\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\alpha\beta\gamma\lambda}$. Thanks to the fact that the surface term $\Theta^{\mu}_{gen}$ is adequately determined by the second term at the right hand side of Eq. (83), it is only necessary to deal with this term regardless of the scalar $\check{S}_{\mu\nu}\delta{g}^{\mu\nu}$ for the sake of deriving the expression for the equations of motion out of the surface term $\Theta^{\mu}_{gen}$. As a consequence of doing so, the scalar $\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}$ is connected with the divergence of the surface term $\Theta^{\mu}_{gen}$ in the way
$\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}=\check{N}_{\mu\nu}\delta{g}^{\mu\nu}+\nabla_{\mu}\Theta^{\mu}_{gen}\,$, (84)
where the rank-two symmetric tensor $\check{N}_{\mu\nu}$ is given by
$\check{N}^{\mu\nu}=\check{B}^{\mu\nu}-\check{P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}-2\nabla_{\rho}\nabla_{\sigma}\check{P}^{\rho\mu\nu\sigma}\,$, (85)
with $\check{B}^{\mu\nu}={B}^{\mu\nu}|_{{Z}\rightarrow\check{Z}}$ and $\check{P}^{\mu\nu\rho\sigma}={P}^{\mu\nu\rho\sigma}|_{{Z}\rightarrow\check{Z}}$, while the surface term $\Theta^{\mu}_{gen}=\Theta^{\mu}|_{{Z}\rightarrow\check{Z}}$ takes the explicit form
$\Theta^{\mu}_{gen}=2\check{P}^{\mu\nu\rho\sigma}\nabla_{\sigma}\delta{g}_{\rho\nu}-2(\delta{g}_{\nu\rho})\nabla_{\sigma}\check{P}^{\mu\nu\rho\sigma}+\sum^{m}_{i=0}\bar{\Theta}^{\mu}_{(i)}\big|_{{Z}\rightarrow\check{Z}}\,$, (86)
with $\bar{\Theta}^{\mu}_{(i)}$ given by Eq. (9). Here it is important to emphasize that equation (84) holds true for any nonnegative integer $m$ and arbitrary tensors $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$s regardless of whether such tensors arise as the results from the variation of the Lagrangian density ${L}_{gen}$ or not.
Moreover, by analogy with the derivation for the identity (133), replacing the variation operator $\delta$ in the scalar $\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}$ with the Lie derivative $\mathcal{L}_{\zeta}$ along the arbitrary smooth vector field $\zeta^{\mu}$ results in the decomposition that this scalar is written as the linear combination of a term proportional to $\mathcal{L}_{\zeta}{g}^{\mu\nu}$ with a divergence term, namely,
$\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}=\check{N}_{\mu\nu}\mathcal{L}_{\zeta}{g}^{\mu\nu}+2\nabla_{\mu}\left(\zeta_{\nu}\check{X}^{\mu\nu}\right)\,.$ (87)
Within the above equation, the rank-two tensor $\check{X}^{\mu\nu}$ is defined as
$\check{X}^{\mu\nu}={X}^{\mu\nu}|_{{Z}\rightarrow\check{Z}}=\check{P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}-2\nabla_{\rho}\nabla_{\sigma}\check{P}^{\rho\mu\nu\sigma}+\check{W}^{\mu\nu}\,.$ (88)
Here $\check{W}^{\mu\nu}=\sum^{m}_{i=0}{W}^{\mu\nu}_{(i)}\big|_{{Z}\rightarrow\check{Z}}$ with the $W^{\mu\nu}_{(i)}$ tensor given by Eq. (23). After a straightforward calculation, the covariant divergence of the $\check{X}^{\mu\nu}$ tensor turns into
$\nabla_{\mu}\check{X}^{\mu\nu}=\frac{1}{2}\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}\nabla^{\nu}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\alpha\beta\rho\sigma}\,.$ (89)
Equation (89) is a necessary condition for the validity of the identity (87). In the same way, the identity (87) holds for any $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$ tensor.
What is more, we observe that the pair $\big(\check{N}^{\mu\nu},\check{X}^{\mu\nu}\big)$ appearing at the right hand side of the identity (87) is non-unique. Here both the rank-two tensors $\check{N}^{\mu\nu}$ and $\check{X}^{\mu\nu}$ denote the symmetric tensor proportional to $\mathcal{L}_{\zeta}{g}^{\mu\nu}$ and the one contracted with $\zeta_{\nu}$ in the divergence term, respectively. For the sake of seeing this clearly, in addition to the pair $\big(\check{N}^{\mu\nu},\check{X}^{\mu\nu}\big)$, it is assumed that there exists another arbitrary pair for rank-two tensors $\big(\check{\mathcal{N}}^{\mu\nu},\check{\mathcal{X}}^{\mu\nu}\big)$ being independent of the vector field $\zeta^{\mu}$, in which the $\check{\mathcal{N}}^{\mu\nu}$ tensor is required additionally to be symmetric. It can be proved that a sufficient and necessary condition for
$\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ $=$ $\check{N}_{\mu\nu}\mathcal{L}_{\zeta}{g}^{\mu\nu}+2\nabla_{\mu}\left(\zeta_{\nu}\check{X}^{\mu\nu}\right)$ (90)
$=$ $\check{\mathcal{N}}_{\mu\nu}\mathcal{L}_{\zeta}{g}^{\mu\nu}+2\nabla_{\mu}\left(\zeta_{\nu}\check{\mathcal{X}}^{\mu\nu}\right)\,$
is that there exists a certain rank-two symmetric and divergence-free tensor $\Omega^{\mu\nu}$ that is irrelevant to the vector field $\zeta^{\mu}$, namely, this tensor has to fulfill
$\Omega^{\mu\nu}=\Omega^{\nu\mu}\,,\qquad\nabla_{\mu}\Omega^{\mu\nu}=0\,$, (91)
rendering the pair $\big(\check{N}^{\mu\nu},\check{X}^{\mu\nu}\big)$ connected with the one $\big(\check{\mathcal{N}}^{\mu\nu},\check{\mathcal{X}}^{\mu\nu}\big)$ in the following manner
$\check{N}^{\mu\nu}=\check{\mathcal{N}}^{\mu\nu}+\Omega^{\mu\nu}\,,\qquad\check{X}^{\mu\nu}=\check{\mathcal{X}}^{\mu\nu}+\Omega^{\mu\nu}\,.$ (92)
As a matter of fact, it is easy to see that there are infinite number of rank-two tensors satisfying the condition (91), which accordingly bring forth an infinite number of pairs $\big(\check{\mathcal{N}}^{\mu\nu},\check{\mathcal{X}}^{\mu\nu}\big)$ fulfilling Eq. (87). However, according to Eq. (92), once either of $\check{\mathcal{N}}^{\mu\nu}$ and $\check{\mathcal{X}}^{\mu\nu}$ is fixed, the other one of them must be determined uniquely. This implies that $\check{\mathcal{N}}^{\mu\nu}$ and $\check{\mathcal{X}}^{\mu\nu}$ are in one-to-one correspondence. Therefore, within Eq. (87), when the symmetric tensor $\check{N}^{\mu\nu}$ is constrained to inherit the one proportional to the variation of the metric in Eq. (84), which is given by Eq. (85), the $\check{X}^{\mu\nu}$ tensor takes the unique form (88). Apparently, owing to the fact that the $\Omega^{\mu\nu}$ tensor is symmetric and divergence-free, it has the potential to convert to the expression for the equations of motion associated to the theories of pure gravity endowed with diffeomorphism invariance if another appropriate pair is supplied in addition to the one $\big(\check{N}^{\mu\nu},\check{X}^{\mu\nu}\big)$. This will be seen below.
In terms of the surface term $\Theta^{\mu}_{gen}$, the straightforward computation demonstrates that the counterpart of this term under the substitution $\delta\rightarrow\mathcal{L}_{\zeta}$ accordingly exhibits the form
$\boxed{\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})=2\zeta_{\nu}\check{X}^{\mu\nu}-\nabla_{\nu}\check{Q}^{\mu\nu}}\,$, (93)
with $\check{Q}^{\mu\nu}={Q}^{\mu\nu}|_{{Z}\rightarrow\check{Z}}$. Like the ${Q}^{\mu\nu}$ tensor, here the anti-symmetric tensor $\check{Q}^{\mu\nu}$ is determined up to the divergence of an arbitrary three-form as well. Equation (93) actually provides a direct way to verify that the vector $\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})$ can be always decomposed into the form (101). Because of the allowance for the arbitrariness of the tensor $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$, the decomposition for $\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})$ displayed by Eq. (93) possesses an evident universality, and it merely depends upon all the $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}$ terms regardless of whether or not such terms stem from the variation of the Lagrangian $\sqrt{-g}{L}_{gen}$. In particular, within the situation where each of the $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}$ terms turns into a component arising from the variation of the Lagrangian, and the sum of all of them fulfills Eq. (83), in the same spirit as the analysis performed towards the derivation of Eq. (141) within Appendix B, it is able to be verified that equation (93) can also appear as a consequence out of the coexistence for both the two equations (83) and (100) with $\delta$ substituted by $\mathcal{L}_{\zeta}$. Additionally, within the same situation, the rank-two tensor $\check{X}^{\mu\nu}$ is symmetric. Nevertheless, the symmetry of this tensor can not be surely guaranteed in other generic cases. What is more, by the aid of Eq. (93), equation (87) is able to be interpreted as the direct consequence from replacing $\delta$ in Eq. (84) with $\mathcal{L}_{\zeta}$, which gives rise to
$\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}=\check{N}_{\mu\nu}\mathcal{L}_{\zeta}{g}^{\mu\nu}+\nabla_{\mu}\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})\,.$ (94)
In other words, for arbitrary tensors $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$s, it is verified via direct calculations that equation (84) with $\delta$ transformed into $\mathcal{L}_{\zeta}$ is valid identically.
On the basis of the surface term (86), making use of the condition that equation (83) still holds after the variation operator $\delta$ is replaced with the Lie derivative $\mathcal{L}_{\zeta}$ along the arbitrary smooth vector field $\zeta^{\mu}$, namely,
$\mathcal{L}_{\zeta}{L}_{gen}=\check{S}_{\mu\nu}\mathcal{L}_{\zeta}{g}^{\mu\nu}+\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}\,$, (95)
we can straightforwardly extend all the related results associated to the Lagrangian (1) to the corresponding ones for the Lagrangian $\sqrt{-g}{L}_{gen}$ under the following substitution
$\frac{\partial{L}}{\partial{g}^{\mu\nu}}\rightarrow\check{S}_{\mu\nu}\,,\qquad\frac{\partial{L}}{\partial\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}}={Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\rightarrow\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\,.$ (96)
Particularly, equation (95) suffices to lead to that the expression ${E}^{\mu\nu}_{gen}$ for the field equations arising from the variation of the Lagrangian (82) can be built from the rank-two tensor $\check{X}^{\mu\nu}$ extracted out of the surface term in the one-to-one manner, that is,
${E}^{\mu\nu}_{gen}=\check{X}^{\mu\nu}-\frac{1}{2}{L}_{gen}{g}^{\mu\nu}\,.$ (97)
Equation (95) guarantees that the $\check{X}^{\mu\nu}$ tensor must be symmetric. In addition, it is worth mentioning that Eq. (95) is a sufficient and necessary condition for Eq. (97) and the conservation law $\nabla_{\mu}{E}^{\mu\nu}_{gen}=0$.
Apart from the aforementioned derivation for the expression (97) relative to the equations of motion, with the help of the Lie derivative of the Lagrangian density along the vector field $\zeta^{\mu}$ presented by Eq. (95), the scalar $\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ is also able to produce such an expression. Concretely, when this scalar is embedded into Eq. (95), yielding another pair
$\big(\check{\mathcal{N}}^{\mu\nu},\check{\mathcal{X}}^{\mu\nu}\big)=\left(\frac{1}{2}{L}_{gen}{g}^{\mu\nu}-\check{S}^{\mu\nu},\quad\frac{1}{2}{L}_{gen}{g}^{\mu\nu}\right)\,$, (98)
by virtue of Eqs. (90) and (92), the comparison between Eqs. (87) and (95) leads to that the rank-two symmetric and divergence-free tensor $\Omega^{\mu\nu}=\check{N}^{\mu\nu}-\check{\mathcal{N}}^{\mu\nu}=\check{X}^{\mu\nu}-\check{\mathcal{X}}^{\mu\nu}$ has to take the specific value
$\Omega^{\mu\nu}$ $=$ $\check{N}^{\mu\nu}+\check{S}^{\mu\nu}-\frac{1}{2}{L}_{gen}{g}^{\mu\nu}$ (99)
$=$ $\check{X}^{\mu\nu}-\frac{1}{2}{L}_{gen}{g}^{\mu\nu}\,.$
Consequently, we obtain the expression for the field equations ${E}^{\mu\nu}_{gen}=\Omega^{\mu\nu}$ together with $\nabla_{\mu}{E}^{\mu\nu}_{gen}=0$ once again. This implies that the $\Omega^{\mu\nu}$ tensor must convert to the expression for the equations of motion if the scalar $\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ is fully determined by the Lie derivative of the Lagrangian density. Notably, here this scalar is required to be put into the form (87) in advance, which reserves the $\check{N}_{\mu\nu}$ tensor in the term $\check{N}_{\mu\nu}\delta{g}^{\mu\nu}$ of Eq. (84) and takes the alternative form (94). Within Eq. (99), the first equality represents nothing else but the expression for the Euler-Lagrange equations of motion, while the second equality denotes the expression of the field equations arising from the surface term. As a result from Eq. (99), it is proved that both of the two expressions must coincide with each other under the condition (95). This further indicates that the symmetric tensor $\check{S}^{\mu\nu}$ is in agreement with the difference between both the tensors $\check{X}^{\mu\nu}$ and $\check{N}^{\mu\nu}$, namely, $\check{S}^{\mu\nu}=\check{X}^{\mu\nu}-\check{N}^{\mu\nu}$, which becomes the identity (43) when the Lagrangian (82) is consistent with the one (1).
Finally, based on the above results in this section, we present the following conclusion:
For any scalar $\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$, where the nonnegative integer $m$ and each of the tensors $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$s are allowed to be arbitrary, the surface term $\Theta^{\mu}_{gen}$ arising from such a scalar is presented by equation (86), which can be always decomposed into the form (93) in terms of the second-rank tensor $\check{X}^{\mu\nu}$ and the anti-symmetric one $\check{Q}^{\mu\nu}$ under the transformation $\delta\rightarrow\mathcal{L}_{\zeta}$. Particularly, when the scalar $\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ is embedded in equation (83) representing the variation of the Lagrangian density ${L}_{gen}$ within equation (82), the $\check{X}^{\mu\nu}$ tensor resulting from the surface term must be able to participate in the construction of the expression ${E}^{\mu\nu}_{gen}$ for the field equations corresponding to the Lagrangian $\sqrt{-g}{L}_{gen}$ in such a way that ${E}^{\mu\nu}_{gen}=\check{X}^{\mu\nu}-\frac{1}{2}{L}_{gen}{g}^{\mu\nu}$ under the condition that equation (95), i.e. equation (83) with $\delta$ substituted by $\mathcal{L}_{\zeta}$, is valid. Besides, the anti-symmetric tensor $\check{Q}^{\mu\nu}$ can be regarded as the Noether charge two-form associated to this Lagrangian.
8 Conclusions and discussions
Within the present paper, starting from the surface term (18) stemming from the variation equation (15) for the generic Lagrangian (1) describing diffeomorphism invariant gravity theories without matter fields, it has been explicitly illustrated that this surface term with the variation operator $\delta$ replaced by the Lie derivative $\mathcal{L}_{\zeta}$ along an arbitrary smooth vector field $\zeta^{\mu}$, denoted by $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$, can be put into the form (101), which is specific to the one given by Eq. (25) for the Lagrangian (1). Such a form for the surface term $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ consists of the symmetric tensor $X^{\mu\nu}$ together with the anti-symmetric one ${Q}^{\mu\nu}$. Subsequently, by utilizing a crucial equation (43), it has been proved that the rank-two symmetric tensor $X^{\mu\nu}$ is in agreement with the functional derivative of the Lagrangian density L{L} with respect to the metric, that is, ${X}_{\mu\nu}=\frac{\delta{L}}{\delta{g}^{\mu\nu}}$. Apparently, as a consequence of this equality, the expression $\tilde{E}^{\mu\nu}={X}^{\mu\nu}-\frac{1}{2}{L}{g}^{\mu\nu}$ resulting from the surface term is completely in agreement with the one ${E}^{\mu\nu}$ for the Euler-Lagrange equations of motion given by Eq. (16). On the other hand, it has been illustrated that the relationship $\tilde{E}^{\mu\nu}={E}^{\mu\nu}$ can be employed to produce the off-shell Noether current ${J}^{\mu}$ in Eq. (49). Its corresponding Noether charge two-form turns out to be the anti-symmetric tensor ${Q}^{\mu\nu}$. Hence we conclude that the equations of motion and the Noether charge can be produced in a uniform framework on the basis of the surface term, that is,
[…]
To see clearly the procedure for the derivation of the field equations, the entire process of doing so is summarized as Eq. (51). In fact, this process can be thought of as the inverse of the one given by Eq. (53), according to which the equations of motion can be also derived out of the surface term as is what has been demonstrated within Ref. [20]. Through the combination of Eq. (51) with (53), the appearance of the form (25) for the surface term, together with either of the two equations (39) and (52), guarantees that there exist some equivalence relationships relative to the expressions for the field equations and the conserved current, exhibited by Eqs. (54) and (56).
In particular, we have illustrated that both the equations (39) and (52), resulting from the substitution of the variation operator by the Lie derivative along the diffeomorphism generator within the variation equations of the Lagrangian in two different forms, are equivalent to each other according to Eqs. (63) and (64) or according to the two equations (141) and (142). Due to this, at the final stage to confirm that both the expressions ${E}^{\mu\nu}$ and $\tilde{E}^{\mu\nu}$ for the equations of motion coincide with each other, it has been demonstrated that each of the equations (39) (or Eq. (43) derived from it) and (52) can serve as the sufficient condition for achieving this. However, by comparison, equation (52) is in advantage by virtue of its conciseness. As a matter of fact, on the basis of Eq. (63), we are able to arrive at the significant conclusion that the validity for either of Eqs. (39) and (52) is a sufficient and necessary condition for that the rank-two tensor ${X}^{\mu\nu}$ can be associated with the Euler-Lagrange expression ${E}^{\mu\nu}$ for the field equations through the one-to-one correspondence ${E}^{\mu\nu}={X}^{\mu\nu}-\frac{1}{2}{L}{g}^{\mu\nu}$ and the covariant divergence for ${E}^{\mu\nu}$ fulfills $\nabla_{\mu}{E}^{\mu\nu}=0$ simultaneously. Inspired with this conclusion, we further provide the generic procedure to derive the expression for the equations of motion from the surface term perspective within Section 7.
Apart from both the two equations (39) and (52), the structure (101) possessed by the surface term with the variation operator $\delta$ transformed into the Lie derivative $\mathcal{L}_{\zeta}$ plays an indispensable role in the construction of the expression for the field equations from the view of the surface term. Due to this, within the framework of any diffeomorphism invariant theory of gravity described by the Lagrangian (82) depending upon the metric, its Riemann tensor and the arbitrary order covariant derivatives of the Riemann tensor, it has been verified through Eq. (93) pertaining to the surface term (86) that the surface term can be always decomposed into the form (101). That is to say, the surface term with $\delta\rightarrow\mathcal{L}_{\zeta}$ must be able to supply the rank-two symmetric tensor that is in agreement with the functional derivative of the Lagrangian density with respect to the metric under the condition (95).
In terms of the results here together with the ones given by our previous work [20], we put forward that the expression for the field equations associated to the diffeomorphism invariant theories of gravity can be extracted out of the surface term. To do so, according to Eq. (50) and the three steps for the derivation of the field equations summarized at the beginning of Section 7, first, varying the Lagrangian density yields the surface term. Next, under the substitution $\delta\rightarrow\mathcal{L}_{\zeta}$, the surface term is written as the form (101). Finally, due to the conclusion that the second-rank tensor contracted with the vector field $\zeta^{\mu}$ in Eq. (101) is identified with the functional derivative of the Lagrangian density $L$ with regard to the metric, the expression for the field equations turns out to be this second-rank tensor minus $\frac{1}{2}{L}{g}^{\mu\nu}$. By contrast with the conventional approach to express the field equations as the Euler-Lagrange equations through the variation of the Lagrangian, resorting to the surface term for producing equations of motion at least has the following merits. First, the derivative of the Lagrangian density with respect to the metric is completely excluded throughout the entire process for the derivation of the field equations. This simplifies the field equations greatly. Second, on the basis of Eq. (56) (see Fig. 1 as well), we find that the surface term provides the advantage of revealing the underlying connections between the conserved current and the expression for the equations of motion by establishing their equivalence relationships in several ways. This straightforwardly supports that the diffeomorphism invariance symmetry of the Lagrangian can determine the equations of motion for the gravitational field in addition to give rise to the conserved current. Third, as what has been shown, the surface term with the variation operator transforming into the Lie derivative can be expressed as the form encompassing a symmetric tensor and an anti-symmetric one, corresponding to the functional derivative of the Lagrangian density with respect to the metric and the Noether charge two-form respectively. Due to this, the surface term offers a unified description for the field equations and the Noether charge. Therefore, apart from the former, the latter can be also derived at the same time. Fourth, the expression for the field equations arising from the surface term provides great convenience for calculating its covariant divergence. Fifth, during the process to derive the field equations, we obtain the explicit expressions for the surface term and the Noether charge two-form simultaneously. These quantities can be directly employed to define the conserved charges of the gravity theories.
Within the context of the pure gravity theories admitting diffeomorphism invariance symmetry, according to Eq. (56) or (113), we obtain the conclusion that the expression for the field equations can be derived out of the vanishing of its covariant divergence. Besides, according to Eq. (56) or (114), the expression for the field equations can also result from the disappearance of the divergence for the off-shell Noether current.
What is more, we have demonstrated that the expression $\tilde{E}^{\mu\nu}$ for the field equations arising from the surface term can be regarded as the most natural generalization of the one associated to the theories of gravity only depending upon the metric and the Riemann tensor given by Ref. [6] to the higher derivative theories of gravity in the presence of the arbitrary order covariant derivatives of the Riemann tensor. To this point, our procedure for the extraction of the field equations out of the surface term offers an alternative path towards the reproduction and the generalization for the field equations through the method proposed by Padmanabhan in Ref. [6]. Inspired with this, we have further revealed that Eq. (61) or the first equivalence relationship in Eq. (63) encompasses the main implications of Padmanabhan’s method. According to Eqs. (50) and (63), although Padmanabhan’s method and our approach from the perspective of surface term produce the same equations of motion, there exist apparent differences between both of them. In fact, within the context of Padmanabhan’s method, the concrete Euler-Lagrange equations of motion arising from the variation of the Lagrangian has to be supplied and it is unavoidable to deal with the Lie derivative of the Lagrangian density to get an equality associated to its derivative with respect to the metric tensor. In contrast, our approach completely abandons the above, and it only involves the handling of the surface term derived out of the variation of the Lagrangian density.
Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant No. 12565009.
Appendix A The structure of the surface term with the variation operator substituted by the Lie derivative
A.1 General formalism
The decomposition of the surface term under the transformation $\delta\rightarrow\mathcal{L}_{\zeta}$ plays a fundamental role in deriving our main results. The desired decomposition form for this quantity gives rise to two tensors corresponding to the expression for equations of motion and the Noether charge two-form respectively. Consequently, for the sake of importance and completeness, here we give some discussions on the structure of the surface term by partially following our earlier work [20].
Without loss of generality, focusing on the generic Lagrangian (82) accommodating diffeomorphism invariance symmetry, that is, $\sqrt{-g}{L}_{gen}=\sqrt{-g}{L}_{gen}\left({g}^{\mu\nu},{R}_{\alpha\beta\rho\sigma},\nabla_{\lambda}\right)$, we vary this Lagrangian with respect to the metric, leading to [1, 21, 22]
$\delta\big(\sqrt{-g}{L}_{gen}\big)=\sqrt{-g}\left(-{E}^{\mu\nu}_{gen}\delta{g}_{\mu\nu}+\nabla_{\mu}\Theta^{\mu}_{gen}\right)\,.$ (100)
As usual, the symmetric rank-two tensor ${E}^{\mu\nu}_{gen}$ in Eq. (100) represents the expression for the Euler-Lagrange equations of motion. For the purpose in this work, as well as in our previous work [20], the surface term $\Theta^{\mu}_{gen}$ with $\delta$ replaced by $\mathcal{L}_{\zeta}$ is supposed to be decomposed into the following general form
$\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})=2\zeta_{\nu}(\text{rank-2 tensor})^{\mu\nu}-\nabla_{\nu}(\text{{anti-symmetric tensor}})^{\mu\nu}\,.$ (101)
[…]
$\nabla_{\mu}{E}^{\mu\nu}_{gen}=0\,.$ (102)
To see this, taking into consideration of the situation where the variation of the Lagrangian is carried out through the diffeomorphism generated by the arbitrary vector field $\zeta^{\mu}$, one is able to replace the variation operator $\delta$ with the Lie derivative $\mathcal{L}_{\zeta}$. Doing so gives rise to the following identity
$\nabla_{\mu}\left[2\zeta_{\nu}\left({E}^{\mu\nu}_{gen}+\frac{1}{2}{L}_{gen}{g}^{\mu\nu}\right)-\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})\right]=2\zeta_{\nu}\nabla_{\mu}{E}^{\mu\nu}_{gen}\,.$ (103)
In accordance with the spirit of the work [20], here we stress that the substitution $\delta\rightarrow\mathcal{L}_{\zeta}$ performed towards the variation equation (100) of the Lagrangian is supposed to be allowed so that Eq. (103) holds identically throughout the present appendix. The identity (103) is instrumental in deriving the main results in this appendix. Owing to the disappearance of the right hand side of Eq. (103), then utilizing Poincare lemma gives rise to
$\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})=2\zeta_{\nu}(\text{{symmetric tensor}})^{\mu\nu}-\nabla_{\nu}(\text{{anti-symmetric tensor}})^{\mu\nu}\,$, (104)
[…]
In light of Eq. (104), we further arrive at the stronger conclusion that the surface term $\Theta^{\mu}_{gen}$ under the replacement $\delta\rightarrow\mathcal{L}_{\zeta}$ can be decomposed into the linear combination for the divergence of a second-rank anti-symmetric tensor together with the contraction between the vector field $\zeta^{\mu}$ and another second-rank symmetric tensor. However, if $\Theta^{\mu}_{gen}$ is supposed to be decomposed as $\Theta^{\mu}_{gen}=\sum_{k=1}\Theta^{\mu}_{(k)}$, it is unnecessary to require each component $\Theta^{\mu}_{(k)}(\delta\rightarrow\mathcal{L}_{\zeta})$ to possess the same structure as the one of Eq. (104).
Apart from the aforementioned consequence that equation (101) can be derived out of the condition $\nabla_{\mu}{E}^{\mu\nu}_{gen}=0$, there exists the fact that its inverse holds true as well, that is, the covariant divergence of the expression for the field equations vanishes identically if the surface term under the substitution $\delta\rightarrow\mathcal{L}_{\zeta}$ can be written as the form (101). To do so, according to Eq. (101), we perform a decomposition of the surface term $\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})$ into two components. One of them consists of the contraction between the vector field $\zeta^{\mu}$ and the rank-two tensor ${X}^{\mu\nu}_{gen}$, which is irrelevant to the former, and the other is the divergence for the anti-symmetric rank-two tensor ${Q}^{\mu\nu}_{gen}$. Specifically, $\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})$ exhibits the form
$\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})=2\zeta_{\nu}{X}^{\mu\nu}_{gen}-\nabla_{\nu}{Q}^{\mu\nu}_{gen}\,.$ (105)
Apparently, by comparison between Eqs. (93) and (105), it is easy to see that both the two rank-two tensors ${X}^{\mu\nu}_{gen}$ and ${Q}^{\mu\nu}_{gen}$ given by the above equation can be specific to ${X}^{\mu\nu}_{gen}=\check{X}^{\mu\nu}$ and ${Q}^{\mu\nu}_{gen}=\check{Q}^{\mu\nu}$, respectively. After substituting Eq. (105) into (103), we have
$\zeta_{\nu}\nabla_{\mu}\left({X}^{\mu\nu}_{gen}-\frac{1}{2}{L}_{gen}{g}^{\mu\nu}\right)+\left({X}^{\mu\nu}_{gen}-{E}^{\mu\nu}_{gen}-\frac{1}{2}{L}_{gen}{g}^{\mu\nu}\right)\nabla_{\mu}\zeta_{\nu}=0\,.$ (106)
Here equation (106) holds for any vector field $\zeta^{\mu}$. Consequently, as is what has been shown in Ref. [20], the expression ${E}^{\mu\nu}_{gen}$ for the equations of motion can be reproduced by means of Eq. (106), that is,
${E}^{\mu\nu}_{gen}={X}^{\mu\nu}_{gen}-\frac{1}{2}{L}_{gen}{g}^{\mu\nu}\,.$ (107)
Apart from this, equation (106) enables one to write down the following identity
$\nabla_{\mu}\left({X}^{\mu\nu}_{gen}-\frac{1}{2}{L}_{gen}{g}^{\mu\nu}\right)=\nabla_{\mu}{E}^{\mu\nu}_{gen}=0\,.$ (108)
Due to the above, one observes that equation (102) is a necessary and sufficient condition for the existence of the form (101) in which the surface term $\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})$ can be, namely,
[…]
Here equation (109) provides another way to verify the disappearance of the covariant divergence for the expression of the field equations, that is, it is only required to confirm that $\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})$ can be decomposed into the form (101). If this fails, the generalized Bianchi identity $\nabla_{\mu}{E}^{\mu\nu}_{gen}=0$ must break down. As a matter of fact, for the diffeomorphism invariant Lagrangian (82), repeating the related analysis for the Lagrangian (1), we find that the surface term $\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})$ is able to be expressed as the form (101). Consequently, $\nabla_{\mu}{E}^{\mu\nu}_{gen}=0$ holds identically. Then identity (103) turns into
$\nabla_{\mu}{J}^{\mu}_{gen}=0\,$, (110)
in which the conserved current ${J}^{\mu}_{gen}$ is defined by
${J}^{\mu}_{gen}=2\zeta_{\nu}{E}^{\mu\nu}_{gen}+\zeta^{\mu}{L}_{gen}-\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})\,.$ (111)
According to the connection between the divergences for ${J}^{\mu}_{gen}$ and ${E}^{\mu\nu}_{gen}$ given by the identity (103), equation (109) is enhanced as
[…]
In addition to this, attributed to the fact that $\nabla_{\mu}{E}^{\mu\nu}_{gen}=0$ guarantees the decomposition for $\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})$ given by Eq. (105) and the ${X}^{\mu\nu}_{gen}$ tensor determined by this equation is in one-to-one correspondence to ${E}^{\mu\nu}_{gen}$ displayed by Eq. (107), one arrives at an interesting conclusion
$\nabla_{\mu}{E}^{\mu\nu}_{gen}=0\quad\Rightarrow\quad{E}^{\mu\nu}_{gen}={X}^{\mu\nu}_{gen}-\frac{1}{2}{L}_{gen}{g}^{\mu\nu}\,.$ (113)
This indicates that the expression for the field equations even can arise from the condition that it is divergence-free. Moreover, equation (112) shows that Eq. (113) is equivalent to
$\nabla_{\mu}{J}^{\mu}_{gen}=0\quad\Rightarrow\quad{E}^{\mu\nu}_{gen}={X}^{\mu\nu}_{gen}-\frac{1}{2}{L}_{gen}{g}^{\mu\nu}\,.$ (114)
That is to say, the expression for the field equations can be also derived out of the divergence-free equation for the conserved current.
With Eq. (105) in hand, we turn our attention to proving that the second-rank tensor contracted with the vector field $\zeta^{\mu}$ is unique as long as the surface term $\Theta^{\mu}_{gen}$ under transformation $\delta\rightarrow\mathcal{L}_{\zeta}$ can be expressed as the form (101). That is to say, once $\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})$ takes on the structure displayed by Eq. (101), the $X^{\mu\nu}_{gen}$ tensor is uniquely determined. To see this, apart from Eq. (105), it is assumed that $\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})$ has another similar form
$\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})=2\zeta_{\nu}\hat{X}^{\mu\nu}_{gen}-\nabla_{\nu}\hat{Q}^{\mu\nu}_{gen}\,.$ (115)
As before, here the second-rank tensor ${X}^{\mu\nu}_{gen}$ bears no relation to the vector field $\zeta^{\mu}$, and $\hat{Q}^{\mu\nu}_{gen}$ represents a certain anti-symmetric tensor. Computing the divergence of the difference between Eqs. (105) and (115), we get
$2\zeta_{\nu}\nabla_{\mu}\Big({X}^{\mu\nu}_{gen}-\hat{X}^{\mu\nu}_{gen}\Big)+2\Big({X}^{\mu\nu}_{gen}-\hat{X}^{\mu\nu}_{gen}\Big)\nabla_{\mu}\zeta_{\nu}=\nabla_{\mu}\nabla_{\nu}\Big({Q}^{\mu\nu}_{gen}-\hat{Q}^{\mu\nu}_{gen}\Big)=0\,.$ (116)
Thus, for the validity of the above equation for any vector field $\zeta^{\mu}$, we further have
$\hat{X}^{\mu\nu}_{gen}={X}^{\mu\nu}_{gen}\,.$ (117)
This implies that the ${X}^{\mu\nu}_{gen}$ tensor is unique. As a result of Eq. (107), it has to be symmetric and takes the specific form
${X}^{\mu\nu}_{gen}={g}^{\mu\rho}{g}^{\nu\rho}\frac{\delta{L}_{gen}}{\delta{g}^{\rho\sigma}}=-\frac{\delta{L}_{gen}}{\delta{g}_{\mu\nu}}\,.$ (118)
By the combination of Eqs. (101) with (105), equation (109) can be further recast into
$\nabla_{\mu}{E}^{\mu\nu}_{gen}=0\Leftrightarrow\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})=-2\zeta_{\nu}\frac{\delta{L}_{gen}}{\delta{g}_{\mu\nu}}-\nabla_{\nu}(\text{{anti-symmetric tensor}})^{\mu\nu}\,.$ (119)
Although it is proved that the $X^{\mu\nu}_{gen}$ tensor is unique, this does not hold true for the anti-symmetric tensor $Q^{\mu\nu}_{gen}$. In fact, the $Q^{\mu\nu}_{gen}$ tensor is determined up to the divergence of an arbitrary three-form. However, the covariant divergence of the rank-two anti-symmetric tensor remains the same, that is,
$\nabla_{\nu}(\text{{anti-symmetric tensor}})^{\mu\nu}=\nabla_{\nu}{Q}^{\mu\nu}_{gen}\,.$ (120)
What is more, due to the fact that the expression for field equations ${E}^{\mu\nu}_{gen}=X^{\mu\nu}_{gen}-\frac{1}{2}{L}_{gen}{g}^{\mu\nu}$, the uniqueness of the $X^{\mu\nu}_{gen}$ tensor further determines that ${E}^{\mu\nu}_{gen}$ is unique as well. As a consequence, one arrives at the conclusion that every Lagrangian corresponds to the unique field equations.
A.2 The application in Lagrangian (1)
Let us pay attention to the specific situation related to the Lagrangian (1). As is shown within Section 2, generally, the surface term $\Theta^{\mu}$ resulting from the variation of the Lagrangian (1) can be split into $\Theta^{\mu}_{P}$ and $\bar{\Theta}^{\mu}_{(i)}$s $(i=1,2,\cdot\cdot\cdot,m)$. If each of them under the transformation $\delta\rightarrow\mathcal{L}_{\zeta}$ is able to be expressed as the form (101), their sum naturally inherits such a form. In fact, for $\Theta^{\mu}_{P}$ coming from the scalar ${P}^{\mu\nu\rho\sigma}\delta{R}_{\mu\nu\rho\sigma}$ in the variation of the Lagrangian, the form (20) for $\Theta^{\mu}_{P}(\delta\rightarrow\mathcal{L}_{\zeta})$ obviously belongs to the one in Eq. (101). On the other hand, for the vector $\bar{\Theta}^{\mu}_{(i)}$, the definition of the Lie derivative for tensors ensures that the surface term $\bar{\Theta}^{\mu}_{(i)}$ with $\delta\rightarrow\mathcal{L}_{\zeta}$ always behaves as
$\bar{\Theta}^{\mu}_{(i)}(\delta\rightarrow\mathcal{L}_{\zeta})$ $=$ $\zeta_{\nu}\Phi^{\mu\nu}_{(i)}+\Psi^{\lambda\mu\nu}_{(i)}\nabla_{\nu}\zeta_{\lambda}$ (121)
[…]
where the tensors $\Phi^{\mu\nu}_{(i)}$ and $\Psi^{\lambda\mu\nu}_{(i)}$ are irrelevant to the vector field $\zeta^{\mu}$. From Eq. (121), one observes that the surface term $\bar{\Theta}^{\mu}_{(i)}(\delta\rightarrow\mathcal{L}_{\zeta})$ can be put into the form (101) if
$\Psi^{\lambda\mu\nu}_{(i)}=\Psi^{\lambda[\mu\nu]}_{(i)}\,$. (122)
Specifically, in terms of Eq. (9), by making use of the identity
$\sum^{i}_{k=1}(-1)^{k-1}\Big(\nabla_{\lambda_{k-1}}\cdot\cdot\cdot\nabla_{\lambda_{1}}{Z}^{\lambda_{1}\cdot\cdot\cdot\lambda_{k-1}\mu\lambda_{k+1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}_{(i)}\Big)\mathcal{L}_{\zeta}\nabla_{\lambda_{k+1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\alpha\beta\rho\sigma}$ (123)
$=$ $\zeta_{\nu}\sum^{i}_{k=1}(-1)^{k-1}\Big(\nabla_{\lambda_{k-1}}\cdot\cdot\cdot\nabla_{\lambda_{1}}Z^{\lambda_{1}\cdot\cdot\cdot\lambda_{k-1}\mu\lambda_{k+1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}_{(i)}\Big)\nabla^{\nu}\nabla_{\lambda_{k+1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\alpha\beta\rho\sigma}$
$-{U}^{\lambda\mu\nu}_{(i)}\nabla_{\nu}\zeta_{\lambda}\,$,
the second-rank tensor $\Phi^{\mu\nu}_{(i)}$ is read off as
$\Phi^{\mu\nu}_{(i)}=\sum^{i}_{k=1}(-1)^{k-1}\Big(\nabla_{\lambda_{k-1}}\cdot\cdot\cdot\nabla_{\lambda_{1}}Z^{\lambda_{1}\cdot\cdot\cdot\lambda_{k-1}\mu\lambda_{k+1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}_{(i)}\Big)\nabla^{\nu}\nabla_{\lambda_{k+1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\alpha\beta\rho\sigma}\,$, (124)
and the third-rank tensor $\Psi^{\lambda\mu\nu}_{(i)}$ has the form
$\Psi^{\lambda\mu\nu}_{(i)}$ $=$ $-{U}^{\lambda\mu\nu}_{(i)}+{U}^{(\nu|\mu|\lambda)}_{(i)}+{U}^{(\nu\lambda)\mu}_{(i)}-{U}^{\mu(\nu\lambda)}_{(i)}$ (125)
[…]
Obviously, $\Psi^{\lambda\mu\nu}_{(i)}=\Psi^{\lambda[\mu\nu]}_{(i)}$ naturally obeys the constraint (122). As a consequence, the vector field $\bar{\Theta}^{\mu}_{(i)}(\delta\rightarrow\mathcal{L}_{\zeta})$ is able to be expressed as the form (101).
Therefore, with the condition (122) satisfied, the surface term $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ can be written as the form (25), in which the tensors $X^{\mu\nu}$ and $Q^{\mu\nu}$ are presented by
[…]
$Q^{\mu\nu}$ $=$ $2{P}^{\mu\nu\rho\sigma}\nabla_{\rho}\zeta_{\sigma}+4\zeta_{\rho}\nabla_{\sigma}{P}^{\mu\nu\rho\sigma}-6{P}^{\mu[\nu\rho\sigma]}\nabla_{\rho}\zeta_{\sigma}-\zeta_{\lambda}\sum^{m}_{i=1}\Psi^{\lambda\mu\nu}_{(i)}\,$, (127)
respectively. Due to the above, in order to guarantee that the surface term $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ possesses the form (101), the sole constraint on this term is to confirm that the third-rank tensor $\Psi^{\lambda\mu\nu}_{(i)}$ fulfills the condition (122).
Appendix B The coexistence of equations (39) and (52) arising from the equivalence between equations (5) and (15)
Both the two equations (5) and (15) are two different but equivalent forms for the variation of the Lagrangian (1). They give rise to Eqs. (39) and (52) under the substitution $\delta\rightarrow\mathcal{L}_{\zeta}$, respectively. In the present appendix, we will demonstrate that Eqs. (39) and (52) always coexist through straightforward calculations.
For convenience, we introduce a scalar V{V} defined through
${V}=\sum^{m}_{i=0}{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}-{P}^{\alpha\beta\rho\sigma}\mathcal{L}_{\zeta}{R}_{\alpha\beta\rho\sigma}+{B}^{\mu\nu}\mathcal{L}_{\zeta}{g}_{\mu\nu}\,.$ (128)
By the aid of the two identities (38) and (73), V{V} is written as
${V}$ $=$ $2{W}^{\mu\nu}\nabla_{\mu}\zeta_{\nu}+\zeta_{\nu}\left(\sum^{m}_{i=0}{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}\nabla^{\nu}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\alpha\beta\rho\sigma}-{P}^{\alpha\beta\rho\sigma}\nabla^{\nu}{R}_{\alpha\beta\rho\sigma}\right)$ (129)
$=$ $2{W}^{\mu\nu}\nabla_{\mu}\zeta_{\nu}+2\zeta_{\nu}\nabla_{\mu}{W}^{\mu\nu}$
$=$ $2\nabla_{\mu}\left(\zeta_{\nu}{W}^{\mu\nu}\right)\,.$
This implies that the scalar V{V} is a total divergence term. Through the combination of Eq. (129) with $2\nabla_{\mu}\big(\zeta_{\nu}{W}^{\mu\nu}\big)=\sum^{m}_{i=0}\nabla_{\mu}\bar{\Theta}^{\mu}_{(i)}(\delta\rightarrow\mathcal{L}_{\zeta})$ stemming from Eq. (22), we gain the following identity
${V}=\nabla_{\mu}\left(\sum^{m}_{i=0}\bar{\Theta}^{\mu}_{(i)}(\delta\rightarrow\mathcal{L}_{\zeta})\right)\,.$ (130)
On the other hand, by utilizing the two identities (68) and (69), we have the identity
${P}^{\mu\nu\rho\sigma}\mathcal{L}_{\zeta}{R}_{\mu\nu\rho\sigma}-\big({P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}+2\nabla_{\rho}\nabla_{\sigma}{P}^{\rho\mu\nu\sigma}\big)\mathcal{L}_{\zeta}{g}_{\mu\nu}$
$=2\nabla_{\mu}\left[\zeta_{\nu}\big({P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}-2\nabla_{\rho}\nabla_{\sigma}{P}^{\rho\mu\nu\sigma}\big)\right]\,.$ (131)
With the help of Eq. (20), the identity (131) is expressed in terms of the covariant divergence of $\Theta^{\mu}_{P}(\delta\rightarrow\mathcal{L}_{\zeta})$ through
$\nabla_{\mu}\Theta^{\mu}_{P}(\delta\rightarrow\mathcal{L}_{\zeta})={P}^{\mu\nu\rho\sigma}\mathcal{L}_{\zeta}{R}_{\mu\nu\rho\sigma}-\big({P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}+2\nabla_{\rho}\nabla_{\sigma}{P}^{\rho\mu\nu\sigma}\big)\mathcal{L}_{\zeta}{g}_{\mu\nu}\,.$ (132)
Subsequently, by means of substituting the identity (131) into the one (129) to eliminate the scalar ${P}^{\mu\nu\rho\sigma}\mathcal{L}_{\zeta}{R}_{\mu\nu\rho\sigma}$ within the latter, we obtain another identity
$\sum^{m}_{i=0}{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}=$
$-\big({B}^{\mu\nu}-{P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}-2\nabla_{\rho}\nabla_{\sigma}{P}^{\rho\mu\nu\sigma}\big)\mathcal{L}_{\zeta}{g}_{\mu\nu}$
$+2\nabla_{\mu}\big[\zeta_{\nu}\big({P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}-2\nabla_{\rho}\nabla_{\sigma}{P}^{\rho\mu\nu\sigma}+{W}^{\mu\nu}\big)\big]\,$, (133)
which is reformulated in terms of both the Euler-Lagrange expression ${E}_{\mu\nu}$ and the surface term $\Theta^{\mu}$ with $\delta\rightarrow\mathcal{L}_{\zeta}$ as
$\sum^{m}_{i=0}{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ $=$ $\left({E}_{\mu\nu}-\frac{\partial{L}}{\partial{g}^{\mu\nu}}+\frac{1}{2}{L}{g}_{\mu\nu}\right)\mathcal{L}_{\zeta}{g}^{\mu\nu}$ (134)
$+\nabla_{\mu}\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})\,.$
Here we point out that Eq. (134) does not appear as a consequence out of the imposition for the constraint that $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ has to be decomposed into the form (25) in advance. It is attributed to the fact that adopting the definition for the Lie derivative to straightforwardly compute the covariant divergence of the surface term $\Theta^{\mu}$ given by Eq. (18) with $\delta\rightarrow\mathcal{L}_{\zeta}$ results in the outcome
$\nabla_{\mu}\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})=2\nabla_{\mu}\big[\zeta_{\nu}\big({P}^{\mu\lambda\rho\sigma}{R}^{\nu}_{~\lambda\rho\sigma}-2\nabla_{\rho}\nabla_{\sigma}{P}^{\rho\mu\nu\sigma}+{W}^{\mu\nu}\big)\big]=2\nabla_{\mu}\big(\zeta_{\nu}{X}^{\mu\nu}\big)\,$ (135)
that the quantity $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ is able to participate in Eq. (134). Furthermore, note that the surface term $\Theta^{\mu}$ completely originates from the sum of the scalar ${Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ with respect to $i$ from 0 to $m$ within Eq. (2). Based on this, such a term, determined only by the variation of the Lagrangian density, can be fully responsible for Eq. (135). Due to this, equation (135) is valid identically within the framework of the theories of gravity described by the Lagrangian (1), even within the case where the tensors ${Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$s are allowed to be arbitrary as shown by Eq. (93). This also indicates that we obtain Eq. (135) regardless of the existence or nonexistence for either of Eqs. (39) and (52), although equation (64) has provided a way to derive Eq. (135) in terms of the coexistence of the two equations. What is more, equation (134) can be seen as a consequence directly arising from the substitution of $\delta$ with $\mathcal{L}_{\zeta}$ in Eq. (136) below.
Actually, on the basis of the fact that the scalar $\sum^{m}_{i=0}{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ originates from the variation of the Lagrangian density L{L}, by utilizing the identity (133), we can offer another explanation for that the coexistence of the two equations (39) and (52) is a sufficient condition for giving rise to Eq. (25) as shown by Eq. (64). For the sake of doing so, direct computation shows that the scalar $\sum^{m}_{i=0}{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ is associated with the covariant divergence of the surface term $\Theta^{\mu}$ in the manner
$\sum^{m}_{i=0}{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}=\left({E}_{\mu\nu}-\frac{\partial{L}}{\partial{g}^{\mu\nu}}+\frac{1}{2}{L}{g}_{\mu\nu}\right)\delta{g}^{\mu\nu}+\nabla_{\mu}\Theta^{\mu}\,.$ (136)
[…]
$\mathcal{L}_{\zeta}\mathcal{L}$ $=$ $\sqrt{-g}\left(\mathcal{L}_{\zeta}{L}-\frac{1}{2}{L}{g}_{\mu\nu}\mathcal{L}_{\zeta}{g}^{\mu\nu}\right)$ (137)
$=$ $\sqrt{-g}\left[\Big(\frac{\partial{L}}{\partial{g}^{\mu\nu}}-\frac{1}{2}{L}{g}_{\mu\nu}\Big)\mathcal{L}_{\zeta}{g}^{\mu\nu}+\sum^{m}_{i=0}{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}\right]\,.\quad$
Substituting the identity (134) into the above equation to cancel out the sum of the scalar ${Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ over $i$ leads to
$\mathcal{L}_{\zeta}\mathcal{L}=\sqrt{-g}\left[{E}_{\mu\nu}\mathcal{L}_{\zeta}{g}^{\mu\nu}+\nabla_{\mu}\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})\right]\,$, (138)
which is just Eq. (15) under the substitution $\delta\rightarrow\mathcal{L}_{\zeta}$}. Conversely, equation (137) can be derived out of Eq. (138) by utilizing the identity (134) once again to get rid of the divergence term $\nabla_{\mu}\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ in Eq. (138). Thus, we observe that the identity (134) can be interpreted as a bridge establishing the equivalence between Eqs. (137) and (138) and equation (135) can arise as the difference between such two equations as illustrated in the above. The invalidity of Eq. (135) implies that this equivalence breaks down. Moreover, we have the significant conclusion that Eq. (135) is definitely valid as long as the variation operator $\delta$ in Eqs. (5) and (15) is allowed to be replaced with the Lie derivative $\mathcal{L}_{\zeta}$.
As a natural consequence of the relation between Eqs. (137) and (138), we arrive at the conclusion that the equivalence relationship between Eqs. (5) and (15), denoted by
[…]
causes their equivalence with the variation operator $\delta$ substituted by the Lie derivative $\mathcal{L}_{\zeta}$, that is,
[…]
Furthermore, attributed to the fact that the two equations (39) and (52) are completely equivalent to Eqs. (5) and (15) under $\delta\rightarrow\mathcal{L}_{\zeta}$, respectively, both of them naturally inherit the equivalence relationship displayed by Eq. (140). Due to this, the existence for either of Eqs. (39) and (52) implies their coexistence, while equation (25) or (135) is a necessary condition for the coexistence of both the two equations.
At the end of this appendix, the relations among the three equations (135), (137) and (138) are summarized as
[…]
By utilizing the equivalence relationships
[…]
equation (141) enables us to arrive at the conclusion that the arbitrary two equations among the three ones (25), (39) and (52) can jointly serve as the sufficient condition for producing the remaining one. Such a conclusion has been presented in terms of the combination of Eqs. (63) and (64) within Section 5.